### v1

# EBC-ZIP: Improving Blockwise Crowd Counting with Zero-Inflated Poisson Regression

Abstract
Density map estimation has become the mainstream paradigm in crowd counting. However, most existing methods overlook the extreme sparsity of ground-truth density maps. In real-world crowd scenes, the vast majority of spatial regions (often over 95%) contain no people, leading to heavily imbalanced count distributions. Ignoring this imbalance can bias models toward overestimating dense regions and underperforming in sparse areas. Furthermore, most loss functions used in density estimation are majorly based on MSE and implicitly assume Gaussian distributions, which are ill-suited for modeling discrete, non-negative count data. In this paper, we propose EBC-ZIP, a crowd counting framework that models the spatial distribution of counts using a Zero-Inflated Poisson (ZIP) regression formulation. Our approach replaces the traditional regression loss with the negative log-likelihood of the ZIP distribution, enabling better handling of zero-heavy distributions while preserving count accuracy. Built upon the recently proposed Enhanced Block Classification (EBC) framework, EBC-ZIP inherits EBC’s advantages in preserving the discreteness of targets and ensuring training stability, while further improving performance through a more principled probabilistic loss. We also evaluate EBC-ZIP with backbones of varying computational complexity to assess its scalability. Extensive experiments on four crowd counting benchmarks demonstrate that EBC-ZIP consistently outperforms EBC and achieves state-of-the-art results. Our code is available at: https://github.com/Yiming-M/EBC-ZIP
Index Terms:
crowd counting, zero-inflated Poisson regression, scalability, blockwise prediction, real-time crowd counting
I Introduction
Crowd counting aims to automatically estimate the number of people present in an image or a video frame. In recent years, it has attracted increasing attention due to its broad range of real-world applications, including public safety management [1], urban planning [2], large-scale event monitoring [3], and intelligent transportation systems [4].
Figure 1: Distribution of blockwise count values (block size = 16) across four crowd counting datasets: ShanghaiTech A & B [5], UCF-QNRF [6], and NWPU-Crowd [7]. The vast majority of blocks have a count of 0, highlighting the extreme sparsity of ground-truth blockwise density maps in these datasets.
While directly regressing the total crowd count from an image inherently discards essential spatial information about where people are located, most modern crowd counting methods [5, 8, 9, 10, 11, 12, 13, 14, 15] address this limitation by predicting density maps instead. However, most of them ignore the extreme sparsity of ground-truth density maps. In real-world crowd scenes, the vast majority of spatial regions contain no people, resulting most values of the ground-truth density maps being zero. If this severe imbalance is not properly addressed, it leads to two significant issues. First, models trained with standard regression losses (e.g., MSE or MAE) become biased toward predicting low or zero values, reducing their sensitivity to sparse yet informative regions. Second, the dominance of zero-valued regions weakens the supervisory signal during training, impeding the model’s ability to learn fine-grained density variations.
To address this issue, the majority of the existing methods [5, 9, 10, 11, 16, 17, 13, 18, 15] adopt blockwise prediction, which aggregates counts over fixed-size patches. Although it can alleviate this issue to some extent, it introduces a trade-off: Small block sizes provide little benefit in sparsity reduction, and large block sizes result in the loss of spatial resolution. As illustrated in Fig. 1, even with a block size of 8, the majority of blocks in most crowd counting datasets still have zero count. There have also been some efforts [19, 20, 18] to address the sparsity by training a segmentation model which aims to separate regions containing people from the background. However, most crowd counting datasets do not provide segmentation masks, they have to rely on Gaussian-smoothed ground-truth density maps to generate pseudo ground-truth segmentation masks. This introduces noise into the pseudo segmentation masks, as the head sizes are not annotated in crowd counting. Guo et al.[18] exploited the box annotations provided by NWPU-Crowd [7] to handle this problem, but compared with point annotations, box annotations are much more expensive to acquire. Besides, these methods introduce an extra task to counting, and how to balance the loss terms in such a multi-task setting poses a new challenge.
On the other hand, most loss functions used by existing crowd counting models are majorly based on blockwise mean squared error (MSE) of density maps. These loss functions, however, are grounded in implicit assumptions that the blockwise count values follow Gaussian distributions. Such assumptions are suboptimal as blockwise count values are inherently discrete, non-negative, and often heavily skewed toward zero. A more suitable statistical model for count data is the Poisson distribution, which captures both the discreteness and non-negativity of counts. Nevertheless, standard Poisson regression struggles with the overwhelming presence of zero counts in crowd counting tasks.
Figure 2: Scalability of the proposed EBC-ZIP framework. The $x$-axis represents the number of model parameters (in millions), and the $y$-axis indicates the Mean Absolute Error (MAE; lower is better) on the validation set of the ShanghaiTech Part B dataset [5]. Each circle represents a specific model, where the radius corresponds to the number of floating point operations (FLOPs) required to process a single $1920\times 1080$ image during inference. EBC-ZIP models (in blue) exhibit a favorable trade-off between model size and accuracy, demonstrating better scalability compared to other lightweight (yellow) and heavyweight (green) models.
To address the above mentioned issues, we propose to adopt a Zero-Inflated Poisson regression (ZIP) model, which explicitly distinguishes between two types of zero counts:
  • Structural zeros, which are hypothesized to be regions that are deterministically zero due to their semantic irrelevance. These include typical background areas such as the sky and buildings, as well as non-head body parts and non-central head regions  [^1] In most crowd counting datasets, only head centers are annotated. As a result, non-head regions like torsos and limbs receive zero count and are treated as structurally zero..
  • Sampling zeros, which refer to head-center regions that nonetheless receive zero count due to the limitations of point-based annotation and annotator subjectivity.
Since there structural and sampling zeros are not annotated separately, our ZIP model learns to distinguish these two types of zeros implicitly by optimizing the negative log-likelihood of the ZIP distribution. Unlike segmentation-based methods [19, 20, 18], our ZIP-based method offers several key advantages. First, we do not require any segmentation masks, and thus, no noise is introduced due to head size uncertainty and we do not need extra work on box annotations. Second, by modeling count as a zero-inflated Poisson distribution over image blocks, our framework captures structural sparsity without assuming the presence of clear object boundaries, making it better suited for highly occluded and densely crowded scenes. We choose to implement our ZIP model within the Enhanced Block Classification (EBC) framework [15], as compared with direct blockwise regression, EBC offers better training stability and preserves the discrete nature of count data. The integration of ZIP further improves this framework by allowing it to capture both structural sparsity and count uncertainty. We also investigate the scalability of EBC-ZIP by testing its performance using backbones with different computational complexities, such as MobileNet [21], MobileCLIP [22], CLIP-ConvNeXt [23, 24]. Fig. 2 highlights the scalability of our EBC-ZIP framework, which outperforms existing models [8, 16, 25, 26, 11, 27, 13] in both accuracy (MAE) and computational cost (FLOPs) across different parameter scales.
In summary, the main contributions of our work are as follows:
  • We propose a novel crowd counting framework EBC-ZIP based on Zero-Inflated Poisson (ZIP) regression, which explicitly models the extreme sparsity of ground-truth count distributions by disentangling structural zeros from sampling uncertainty. Built on top of the EBC framework, EBC-ZIP significantly improves its performance by replacing heuristic classification with a principled probabilistic formulation.
  • We present the first systematic study of framework-level scalability in crowd counting. Our EBC-ZIP framework generalizes well across a wide range of model sizes and architectures, from lightweight convolutional neural networks (CNNs) to vision-language models, achieving state-of-the-art performance under varying computational constraints.
  • We conduct extensive experiments on four crowd counting benchmarks to validate the effectiveness of our EBC-ZIP framework. Results show that our base model, EBC-ZIP-Base, consistently outperforms existing state-of-the-art methods across four different datasets.
II Related Work
II-A Handling Sparsity
To address the extreme sparsity of ground-truth density maps, many existing methods [5, 9, 10, 16, 17, 13, 18] adopt a combination of Gaussian smoothing and blockwise prediction. These approaches typically preprocess the binary annotation map by convolving it with Gaussian kernels, either of fixed or adaptive size. Specifically, several works [9, 10, 17, 13, 18] use fixed kernel bandwidths, while others [5, 16] adjust the kernel size according to the distance to the nearest annotated head. In principle, the kernel size should reflect the actual head size in the image; however, such information is typically unavailable in most crowd counting datasets. As a result, Gaussian smoothing inevitably introduces noise into the ground-truth density maps, which can degrade the model’s generalization ability [11]. To overcome these limitations, recent works [11, 15] propose to eliminate Gaussian smoothing altogether and instead rely solely on blockwise prediction. Empirically, these smoothing-free blockwise methods have demonstrated better performances compared to those using Gaussian-smoothed density maps. Nevertheless, the sparsity issue persists: as shown in Fig. 1, even with a block size of 8, which is used in the aforementioned blockwise prediction methods, over 85% of the blocks still have zero count in most crowd counting datasets. This extreme imbalance presents significant challenges for training and optimization, as the vast majority of the model’s supervision signal comes from zeros.
II-B Modeling Count Distributions in Density Maps
Most density-based crowd counting methods typically frame the task as a regression problem, where the model is trained to minimize the blockwise mean squared error (MSE) between the predicted and ground-truth blockwise density maps [5, 9, 28, 10, 16, 13, 18]. This loss function implicitly assumes that each blockwise count follows a Gaussian distribution centered at the predicted value. Although this formulation is simple and effective in many scenarios, it fails to capture the fundamental characteristics of count data, which is discrete, non-negative, and often exhibits a long-tailed distribution. DMCount [11] addresses a different issue by reformulating density estimation as an Earth Mover’s Distance (EMD) problem. By minimizing the Wasserstein distance between predicted and ground-truth density maps, DMCount explicitly models the global transportation cost needed to match the predicted mass to the annotated points. This approach improves robustness to spatial misalignment and annotation noise. However, similar to MSE-based methods, it still operates entirely in a continuous density space, treating the density maps as divisible mass rather than integer-valued counts. Also, the exact computation of EMD scales poorly with the number of bins ($\mathcal{O}(n^{3})$). In contrast, Bayesian Poisson Regression (BPR) is [29] an early attempt to model count from a discrete perspective. In BPR, an image is segmented into large regions, each assumed to follow a Poisson distribution, with the log-mean regressed from hand-crafted low-level features. While this formulation respects the discrete and non-negative nature of count data, the spatial segmentation is often coarse, and as a result, the independence assumption between regions is unrealistic in densely crowded scenes, where spatial correlation is strong. This issue can be mitigated by adopting a finer-grained modeling strategy by treating small blocks (e.g., $8\times 8$ or $16\times 16$) as the basic counting units. Local interactions tend to be weaker under this circumstance, so the assumption of blockwise independence becomes more plausible.
In this work, we propose to model blockwise count distributions using zero-inflated Poisson (ZIP) regression. Unlike existing methods, the zero-inflated Poisson (ZIP) distribution better aligns with the empirical distribution of blockwise crowd counts, as it naturally captures their discreteness, non-negativity, long-tailedness, and zero-heaviness.
III Method
Figure 3: Overview of the proposed EBC-ZIP framework. Given an input image $\boldsymbol{X}\in\mathbb{R}^{3\times H\times W}$, a backbone network (e.g., MobileNet [21], MobileCLIP [22], or CLIP [30]) extracts feature maps, which are shared by two parallel branches. In the EBC Poisson branch (top), the features are passed to a $\boldsymbol{\lambda}$-head with a softmax activation, producing the blockwise count distribution $\boldsymbol{P}^{*}_{\boldsymbol{\lambda}}$ over $n-1$ positive bins. The Poisson rate per block is then computed as a weighted average of bin centers, yielding the final $\boldsymbol{\lambda}\in\mathbb{R}^{1\times h\times w}$ map. This branch is supervised via the cross-entropy loss $\mathcal{L}_{\mathrm{CE}}$. In the Zero Inflation branch (bottom), the same features are fed into a $\boldsymbol{\pi}$-head with a sigmoid activation to estimate the structural zero probability map $\boldsymbol{\pi}\in\mathbb{R}^{1\times h\times w}$. The predicted density map $\boldsymbol{Y}^{*}$ is defined as the expected value of the zero-inflated Poisson distribution: $\boldsymbol{Y}^{*}=(1-\boldsymbol{\pi})\otimes\boldsymbol{\lambda}$, where $\otimes$ denotes elementwise multiplication. The overall model is trained using the ZIP-based negative log-likelihood loss $\mathcal{L}_{\mathrm{NLL}}$ and an auxiliary count loss $\mathcal{L}_{\mathrm{count}}$ to ensure accurate total count prediction.
Our proposed framework, EBC-ZIP, is an improved extension of the Enhanced Block Classification (EBC) framework introduced in our prior work [15]. In this work, we propose to incorporate Zero-Inflated Poisson (ZIP) regression to address the extreme sparsity of ground-truth density maps. We hypothesize that the zeros in the ground-truth can be categorized into two types: structural zeros, which correspond to semantically irrelevant regions, and sampling zeros, which arise from point-based annotation uncertainty. ZIP regression provides a principled way to model this dual-source sparsity and aligns better with the discrete, non-negative nature of count data. For completeness, we first provide a brief overview of the original EBC framework, followed by a detailed description of ZIP regression. We then present our full EBC-ZIP architecture, which combines the benefits of blockwise classification and probabilistic count modeling. An overview of the proposed framework EBC-ZIP is shown in Fig. 3.
III-A Recap: EBC
Training crowd counting models directly via regression suffers from several limitations. For instance, regression models must learn to accurately predict rare but high-impact values, which can lead to unstable gradient updates – especially in the presence of outliers within a batch. Moreover, regression outputs are typically unbounded (e.g., from a linear layer), which can result in negative or extremely large predictions during early training, causing gradient explosion and slow convergence. The Enhanced Block Classification (EBC) framework addresses these issues by reformulating crowd counting as a blockwise classification problem.
In EBC, blockwise count values are quantized into integer-valued bins $\{\mathcal{B}_{k}\}_{k=1}^{n}$, with larger counts grouped into a shared bin to reduce label sparsity and mitigate the effect of annotation errors. For example, in the case of block size $r=8$, EBC utilizes 00{0}, 11{1}1, 22{2}2, 33{3}3, and $[4,\infty)$ as bins.
Given an input image $\boldsymbol{X}\in\mathbb{R}^{3\times H\times W}$, a backbone network $\mathcal{F}$ (e.g., ResNet [31], VGG [32]) is used to extract feature maps:
$\boldsymbol{f}=\mathcal{F}(\boldsymbol{X})\in\mathbb{R}^{C\times h\times w}$ (1)
where $C$ is the number of channels, and $h=H//r$, $w=W//r$ correspond to the blockwise spatial resolution given block size $r$.
A $1\times 1$ convolutional layer $\mathcal{H}$ followed by softmax is applied to produce the predicted blockwise classification probability map:
$\boldsymbol{P}^{*}=\mathrm{Softmax}\left(\mathcal{H}(\boldsymbol{f})\right)\in\mathbb{R}^{n\times h\times w}.$
Here, the vector $\boldsymbol{P}^{*}_{:,i,j}\in\mathbb{R}^{n}$ denotes the categorical distribution over the $n$ predefined bins for block $(i,j)$.
The final predicted blockwise density map $\boldsymbol{Y}^{*}$ is obtained by taking the expectation over bin centers ${b_{k}}_{k=1}^{n}$:
$\boldsymbol{Y}^{*}_{i,j}=\sum_{k=1}^{n}\boldsymbol{P}^{*}_{k,i,j}\cdot b_{k}.$ (2)
While EBC alleviates the challenges of regression through classification, it still does not fully address the extreme sparsity of ground-truth density maps, where most blocks have zero count (see Fig. 1). Also, its DMCount-based loss function contradicts the discrete nature of blockwise counts. To handle these issues, we introduce zero-inflated Poisson (ZIP) regression, which forms the basis of our proposed extension, EBC-ZIP.
Figure 4: Illustration of structural and sampling zeros (best viewed in color). The image (synthesized for illustration purposes) is overlaid with a grid where each cell represents a spatial block. The red dot marks a ground-truth head annotation, and the block containing it is labeled with the red number “1”. Yellow zeros, which indicate sampling zeros, are assigned to blocks that correspond to the head-center region but receive zero count due to the point-based annotation protocol and randomness. Green zeros denote structural zeros, corresponding to background regions, non-head body parts or outer regions of the head that are not associated with any annotations.
(a) MAE
(b) RMSE
(c) NAE (%)
Figure 5: Performance of the VGG19-based structure on ShanghaiTech B under varying values of $\omega$ in the loss function (15). Results indicate that $\omega=0.75$ yields the best MAE (6.34) and RMSE (9.79), while $\omega=1.00$ gives the lowest NAE (5.00%).
III-B Zero-Inflated Poisson Regression
Poisson regression is a classical method for modeling count data, where the observed variable $u\in\mathbb{N}_{+}$ is assumed to follow a Poisson distribution with rate parameter $\lambda>0$. The probability mass function (p.m.f.) is defined as:
$P(u\mid\lambda)=\frac{\lambda^{u}e^{-\lambda}}{u!},\quad u\in\{0,1,2,\cdots\}$ (3)
The Poisson distribution is widely used to model the number of discrete events occurring within a fixed interval of time or space, such as the the number of cars passing through a checkpoint per minute, or the number of phone calls received by a call center in an hour. In the context of crowd counting, it is natural to model the number of people appearing in a spatial region (e.g., an image block) as a Poisson random variable, where $\lambda$ denotes the expected count. To reasonably assume independence between blocks, the block size must be sufficiently small, since person-to-person interactions tend to be local. In contrast, using large regions (such as in early works like [29]) may violate this assumption, as spatial correlations between regions may increase in such a scenario. Consequently, it is inappropriate to represent the overall likelihood as the product of probability mass functions.
However, standard Poisson distribution underestimates the probability of observing zeros when the data is zero-heavy. To explicitly handle this, we adopt the Zero-Inflated Poisson (ZIP) model, which which introduces a latent binary process to account for structural sparsity. The ZIP distribution is a mixture of a point mass at zero and a standard Poisson distribution:
$P(v\mid\pi,\lambda)=\begin{cases}\pi+(1-\pi)e^{-\lambda},\;&v=0\\ (1-\pi)\frac{\lambda^{v}e^{-\lambda}}{v!},\;&v>0\end{cases}$ (4)
Here, $\pi\in[0,1]$ represents the probability of a structural zero, and $\lambda$ is the rate of the Poisson component responsible for modeling counts. The probability of a sampling zero is given by $(1-\pi)\exp({-\lambda})$.
This decomposition aligns well with the inherent structure of crowd counting data. Most zero-valued blocks in ground-truth density maps correspond to semantically irrelevant regions such as background (e.g., sky, buildings), non-head body parts (e.g., torsos, limbs), or peripheral head regions, which are never annotated. We refer to these as structural zeros, and explicitly modeling them helps the model suppress spurious count predictions in these regions. In contrast, some zero-valued blocks occur in semantically valid regions (e.g., the head-center area of a person) but do not receive annotation due to the limitations of point-based labeling: Since each person is annotated using a single point (typically the head center), and a head may cover multiple blocks, only one of these blocks receives a non-zero label, while others remain zero. Additionally, annotator subjectivity can cause annotation points to fall in different blocks across annotators. These blocks are not truly empty but are instead artifacts of annotation sparsity and subjectivity, and are therefore better modeled as sampling zeros, arising from the Poisson component. Fig. 4 illustrates these two types of zeros.
The expectation of $v\sim\mathrm{ZIP}(\pi,\lambda)$ can be derived as:
$\mathbb{E}[v\mid\pi,\lambda]$ $=\sum_{k=0}^{\infty}k\cdot P(v=k\mid\pi,\lambda)$
$=\sum_{k=1}^{\infty}k\cdot P(v=k\mid\pi,\lambda)$
$=(1-\pi)\sum_{k=0}^{\infty}k\cdot\frac{\lambda^{k}e^{-\lambda}}{k!}$
$=(1-\pi)\mathbb{E}[u\mid\lambda]=(1-\pi)\lambda,$ (5)
where $u\sim\mathrm{Poisson}(\lambda)$ and $\mathbb{E}(u\mid\lambda)=\lambda$. This expectation confirms that the predicted count is suppressed proportionally to the probability of structural zeros, yielding a more faithful modeling of blockwise density in sparse crowd scenes.
III-C Extending EBC to EBC-ZIP
Similar to EBC, our proposed EBC-ZIP framework is backbone-agnostic, allowing it to be integrated with various feature extractors. Given an input image $\boldsymbol{X}\in\mathbb{R}^{3\times H\times W}$, we first compute the feature map $\boldsymbol{f}$ using a shared backbone, as described in Eq. (1). We then adopt the EBC framework to predict Poisson rates, due to its training stability and robustness to label errors. Specifically, the feature map $\boldsymbol{f}$ is passed through a $\boldsymbol{\lambda}$-head $\mathcal{H}_{\boldsymbol{\lambda}}$ (implemented as a $1\times 1$ convolution), followed by a softmax activation to generate a probability distribution over the $n{-}1$ positive count bins:
$\boldsymbol{P}^{*}_{\boldsymbol{\lambda}}=\mathrm{Softmax}\left(\mathcal{H}_{\boldsymbol{\lambda}}(\boldsymbol{f})\right)\in\mathbb{R}^{(n-1)\times h\times w}$ (6)
where $h=H//r$, $w=W//r$ denote the height and width of the blockwise output map. Note that we exclude the $\{0\}$ bin here, since the Poisson distribution assumes a strictly positive rate $\lambda>0$. We then compute the estimated Poisson rate map $\boldsymbol{\lambda}\in\mathbb{R}_{+}^{1\times h\times w}$ as the expected value of the bin centers, analogous to Eq. (2):
$\boldsymbol{\lambda}_{i,j}=\sum_{k=1}^{n-1}{\boldsymbol{P}^{*}_{\boldsymbol{\lambda}}}_{k,i,j}\cdot{b_{+}}_{k},$ (7)
where ${b_{+}}_{k}$ for $k=1,\cdots,n-1$ denotes the center of each positive bin.
The learning of $\boldsymbol{P}^{*}_{\boldsymbol{\lambda}}$ is supervised by a cross-entropy loss over the positive regions of the ground-truth density map. Specifically, we define the set of positive-valued blocks as $\boldsymbol{Y}_{+}\coloneqq\boldsymbol{Y}[\boldsymbol{Y}>0]$, which arise from the Poisson component and are thus only related to $\boldsymbol{\lambda}$. Their corresponding probabilistic predictions $\boldsymbol{P}^{*}_{\boldsymbol{\lambda}+}$ can be obtained via $\boldsymbol{P}^{*}_{\boldsymbol{\lambda}+}=\boldsymbol{P}^{*}_{\boldsymbol{\lambda}}[:,\boldsymbol{Y}>0]$ The cross-entropy loss is calculated based on $\boldsymbol{P}^{*}_{\boldsymbol{\lambda}+}$ and $\boldsymbol{Y}_{+}$:
$\mathcal{L}_{\mathrm{CE}}=\mathrm{CrossEntropy}(\boldsymbol{P}^{*}_{\boldsymbol{\lambda}+},\boldsymbol{P}_{+}),$ (8)
where $\boldsymbol{P}_{+}$ is the one-hot encoded ground-truth probability map of $\boldsymbol{Y}_{+}$.
In parallel, we introduce a $\boldsymbol{\pi}$-head, $\mathcal{H}_{\boldsymbol{\pi}}$, also implemented as a $1\times 1$ convolution, followed by a sigmoid activation to produce the structural zero probability map:
$\boldsymbol{\pi}=\sigma\left(\mathcal{H}_{\boldsymbol{\pi}}(\boldsymbol{f})\right)\in\mathbb{R}^{1\times h\times w}$ (9)
Suffix Backbone Size (M) FLOPS (G)
-P (Pico) MobileNetV4-Small-0.5×0.5$\times$0.5 × 0.81 6.46
-N (Nano) MobileNetV4-Small 3.36 24.73
-T (Tiny) MobileNetV4-Medium 10.64 61.62
-S (Small) MobileCLIP-S1 33.72 242.68
-B (Base) CLIP-ConvNeXt-Base 105.60 800.99
TABLE I: EBC-ZIP model variants with corresponding backbones, parameter sizes, and computational complexity measured on $1920\times 1080$ resolution.
Under the assumption that blockwise counts are conditionally independent, the negative log-likelihood (NLL) loss of the zero-inflated Poisson distribution is given by:
$\mathcal{L}_{\mathrm{NLL}}=$ $-\log\prod_{i=1}^{h}\prod_{j=1}^{w}P(\boldsymbol{V}_{i,j}\mid\boldsymbol{\pi}_{i,j},\boldsymbol{\lambda}_{i,j})$
$=$ $-\sum_{i=1}^{h}\sum_{j=1}^{w}\log P(\boldsymbol{V}_{i,j}\mid\boldsymbol{\pi}_{i,j},\boldsymbol{\lambda}_{i,j})$ (10)
where $P(\boldsymbol{V}_{i,j}\mid\boldsymbol{\pi}_{i,j},\boldsymbol{\lambda}_{i,j})$ is defined by the ZIP distribution in Eq. (4), and $\boldsymbol{\lambda}$ and $\boldsymbol{\pi}$ are given by (7) and (9), respectively. It is important to note that we do not supervise the $\boldsymbol{\pi}$ and $\boldsymbol{\lambda}$ branches separately. Since structural and sampling zeros are not explicitly annotated in the ground-truth, we treat them as latent factors and optimize both branches jointly via the ZIP negative log-likelihood loss Eq. (10). This formulation allows the model to implicitly learn to disentangle the two types of zeros based on the statistical patterns in the data, without requiring additional annotations.
We use the expectation map $\boldsymbol{Y}^{*}=\mathbb{E}[\boldsymbol{V}|\boldsymbol{\pi},\boldsymbol{\lambda}]$ as the predicted density map, given by
$\boldsymbol{Y}^{*}_{i,j}=\mathbb{E}[\boldsymbol{V}_{i,j}\mid\boldsymbol{\pi}_{i,j},\boldsymbol{\lambda}_{i,j}].$ (11)
Using Eq. (5), $\boldsymbol{Y}^{*}$ can be expressed as
$\boldsymbol{Y}^{*}=(1-\boldsymbol{\pi})\otimes\boldsymbol{\lambda},$ (12)
where $\otimes$ represents element-wise multiplication.
To ensure accurate crowd estimates at the image level, we incorporate a count loss that penalizes discrepancies between the predicted and ground-truth total counts. The predicted total count $c^{*}$ is obtained by summing over all elements of the predicted density map:
$c^{*}=\sum_{i=1}^{h}\sum_{j=1}^{w}\boldsymbol{Y}^{*}_{i,j}.$ (13)
We define the count loss as the mean absolute error (MAE) between the predicted and ground-truth total counts:
$\mathcal{L}_{\mathrm{count}}=\left|c^{*}-c\right|=\left|\sum_{i=1}^{h}\sum_{j=1}^{w}\left(\boldsymbol{Y}^{*}_{i,j}-\boldsymbol{Y}_{i,j}\right)\right|$ (14)
Finally, the model is trained using a weighted sum of three loss terms:
$\mathcal{L}_{\mathrm{total}}=\omega\mathcal{L}_{\mathrm{CE}}+\mathcal{L}_{\mathrm{NLL}}+\mathcal{L}_{\mathrm{count}},$ (15)
where $\omega$ is a scalar weighting coefficients, and $\mathcal{L}_{\mathrm{CE}}$, $\mathcal{L}_{\mathrm{NLL}}$, and $\mathcal{L}_{\mathrm{count}}$ are defined in Eq. (8), (10), and (14), respectively.
Variants. To evaluate the scalability and adaptability of our proposed framework, we construct five model variants of EBC-ZIP with different backbones, including MobileNetV4 [21], MobileCLIP [22] and OpenCLIP’s ConvNeXt-Base [24, 23] These variants are suitable for different application scenarios, ranging from mobile-friendly to high-performance settings. The specifications are summarized in Table I. All variants share the same head design and differ only in backbone complexity. This modularity allows EBC-ZIP to scale across computational budgets.
Regression Component MAE RMSE NAE
MAE 6.69 10.45 5.52%
DMCount [11] 6.62 9.83 5.30%
Poisson NLL 6.57 9.86 5.47%
ZIP NLL (ours) 6.36 10.32 5.00%
TABLE II: Comparison of ZIP NLL with other choices in the total loss function Eq. (15) on ShanghaiTech B. Other loss terms (i.e., $\mathcal{L}_{\mathrm{CE}}$ and $\mathcal{L}_{\mathrm{count}}$) remain unchanged. Results show that our ZIP NLL achieves the lowest MAE and NAE.
Method ShanghaiTech A ShanghaiTech B UCF-QNRF NWPU-Crowd (Val)
MAE RMSE MAE RMSE MAE RMSE MAE RMSE
BL [10] 62.8 101.8 7.7 12.7 88.7 154.8 93.6 470.3
DMCount [11] 59.7 95.7 7.4 11.8 85.6 148.3 70.5 357.6
EBC [15] 62.3 98.9 7.0 10.9 77.2 130.4 39.6 95.8
EBC-ZIP (ours) 58.6 96.0 6.3 10.3 73.0 126.1 38.3 90.3
Improvement over EBC 5.9% 2.9% 10.0% 5.5% 5.4% 3.2% 3.2% 5.7%
TABLE III: Comparison of EBC-ZIP with other frameworks on ShanghaiTech A & B, UCF-QNRF and NWPU-Crowd. All frameworks use the VGG19-based backbone. Results demonstrate that our framework EBC-ZIP significantly outperforms other methods, achieving the lowest MAE and RMSE on all four benchmarks. Our EBC-ZIP also consistently improves the EBC framework, at most by 10.0%. The best results are highlighted in bold font, and the second best results are underscored.
IV Experiments
Datasets. Following [15], we evaluate all EBC-ZIP variants listed in Table I on four widely used crowd counting benchmarks: ShanghaiTech A & B [5], UCF-QNRF [6], and NWPU-Crowd [7]. ShanghaiTech A contains 300 training images and 182 test images, with highly congested scenes and an average of 501 people per image. ShanghaiTech B, the only dataset collected from surveillance viewpoints, includes 400 training and 316 test images, featuring relatively sparse crowds. UCF-QNRF comprises 1,201 training images and 334 test images with extremely dense crowds, averaging 815 people per image. NWPU-Crowd is currently the largest high-resolution crowd counting dataset, containing 3,109 training images, 500 validation images, and 1,500 test images, with an average count of 418. The test set annotations of NWPU-Crowd are not publicly released. To obtain evaluation results on this split, predicted counts must be submitted to the official evaluation server. This protocol helps maintain a fair comparison between methods by preventing result tweaking using test labels.
Setup. Our training configuration largely follows [15]. We first preprocess all datasets so that the minimum edge length is no less than 448. We limit the maximum edge length of UCF-QNRF to be 2048 and that of NWPU-Crowd to be 3072. On ShanghaiTech A & B and NWPU-Crowd, we use a block size of 16 while on UCF-QNRF we use a block size of 32. We apply a combination of data augmentation techniques, including random resized cropping, horizontal flipping, color jittering, and salt-and-pepper noise injection. For ShanghaiTech A and B, which consist of lower-resolution images, the crop size is set to $448\times 448$. For the higher-resolution datasets UCF-QNRF and NWPU-Crowd, we use a larger crop size of $672\times 672$. All models are optimized using the Adam optimizer with a weight decay of $41e-$ 4. The learning rate is linearly warmed up from $51e-$ 5 to $41e-$ 4 over the first 25 epochs, followed by cosine annealing with parameters $T_{0}=5$ and $T_{\mathrm{mult}}=2$. Each model variant is trained on each dataset for a total of 1,300 epochs with a batch size of 8. All experiments are conducted on a single NVIDIA RTX 5000 GPU.
Evaluation metrics. Following standard practice, we evaluate each model variant on all datasets using three commonly adopted metrics: Mean Absolute Error (MAE), Root Mean Squared Error (RMSE), and Normalized Absolute Error (NAE). The NAE metric ensures fair evaluation across both sparse and dense scenes by normalizing the absolute error with respect to the ground-truth count. These metrics are defined as follows:
$\mathrm{MAE}=\frac{1}{M}\sum_{m=1}^{M}\left|c_{m}-c_{m}^{*}\right|,$ (16)
$\mathrm{RMSE}=\sqrt{\frac{1}{M}\sum_{m=1}^{M}\left|c_{m}-c_{m}^{*}\right|^{2}},$ (17)
and
$\mathrm{NAE}=\frac{1}{\|\mathcal{M}_{+}\|}\sum_{m\in\mathcal{M}_{+}}\frac{\left|c_{m}-c_{m}^{*}\right|}{c_{m}},$ (18)
where $M$ denotes the total number of test images, $c_{m}$ is the ground-truth count for image $m$, and $c_{m}^{*}$ is the predicted count, computed via Eq. (13). The set $\mathcal{M}+\subseteq{1,2,\dots,M}$ includes only the indices for which $c_{m}>0$, and $|\mathcal{M}+|$ denotes its cardinality. The NAE metric excludes empty scenes to avoid division by zero and to better reflect relative error in populated scenes.
IV-A Effectiveness of EBC-ZIP
In this subsection, we empirically validate the effectiveness of our proposed EBC-ZIP framework by comparing it against several strong baselines, including BL [10], DMCount [11], and EBC [15]. For a fair comparison, we adopt a consistent encoder-decoder architecture based on VGG19 [32]. Before proceeding to the comparison, we first determine the optimal value of the weighting parameter $\omega$ in Eq.(15), which balances the cross-entropy term and other two terms. To this end, we evaluate the VGG19-based model under various settings of $\omega$ on the ShanghaiTech B dataset and report the results in Fig.5. As illustrated, the model achieves the best MAE and RMSE when $\omega=0.75$, and the lowest NAE when $\omega=1.00$. In our subsequent comparisons, we fix $\omega=1.00$ to prioritize NAE, due to its robustness to varying crowd densities.
To validate the effectiveness of the proposed ZIP NLL loss, we conduct an ablation study where we replace the ZIP NLL component in the total loss function (15) with several commonly used alternatives, including MAE, DMCount [11], and Poisson NLL. Other components in the loss function, such as the cross-entropy loss and the count loss, are kept unchanged. All models are trained using the same VGG19-based architecture under identical settings. The evaluation results on ShanghaiTech B are presented in Table II. As shown, our ZIP NLL achieves the lowest MAE (6.36) and NAE (5.00%), and performs competitively in RMSE (10.32), demonstrating its effectiveness in enhancing counting accuracy while maintaining robust distribution modeling.
Finally, we compare EBC-ZIP framework with several existing methods on the four benchmarks, shown in Table III. Our approach sets a new state-of-the-art on ShanghaiTech B, UCF-QNRF and NWPU-Crowd. It also achieves the lowest MAE and the second lowest RMSE on ShanghaiTech A. Notably, when compared directly with its EBC baseline, our EBC-ZIP model demonstrates significant performance gains across all datasets, reducing the MAE by a margin of 3.2% to 10.0%. hese improvements validate the efficacy of incorporating zero-inflated Poisson regression for crowd counting.
Method Venue ShanghaiTech A ShanghaiTech B UCF-QNRF NWPU (Val)
MAE RMSE MAE RMSE MAE RMSE MAE RMSE
ChfL [33] CVPR’22 57.5 94.3 6.9 11.0 80.3 137.6 76.8 343.0
MAN [34] CVPR’22 56.8 90.3 - - 77.3 131.5 76.5 323.0
CLTR [35] ECCV’22 56.9 95.2 6.5 10.6 85.8 141.3 61.9 246.3
CrowdHat [36] CVPR’23 51.2 81.9 5.7 9.4 75.1 126.7 - -
STEERER [13] ICCV’23 54.5 86.9 5.8 8.5 74.3 128.3 54.3 238.3
PET [37] ICCV’23 49.3 78.8 6.2 9.7 79.5 144.3 58.5 238.0
APGCC [38] ECCV’24 48.8 76.7 5.6 8.7 80.1 133.6 - -
Gramformer [39] AAAI’24 54.7 87.1 - - 76.7 129.5 - -
mPrompt [18] CVPR’24 52.5 88.9 5.8 9.6 72.2 133.1 50.2 219.0
CLIP-EBC [15] ICME’25 52.5 83.2 6.0 10.1 80.3 136.6 36.6 81.7
EBC-ZIP-B (ours) - 47.81 75.04 5.51 8.63 69.46 121.88 28.26 64.84
TABLE IV: Comparison of our model EBC-ZIP-B with state-of-the-art approaches on ShanghaiTech A & B, UCF-QNRF, and the NWPU-Crowd validation split. EBC-ZIP-B achieves the best performance across all these four datasets under both MAE and RMSE, surpassing all strong baselines.
Method Venue MAE RMSE NAE
ChfL [33] CVPR’22 76.8 343.0 0.171
MAN [34] CVPR’22 76.5 323.0 0.170
CLTR [35] ECCV’22 74.4 333.8 0.165
CrowdHat [36] CVPR’23 68.7 296.9 0.181
STEERER [13] ICCV’23 63.7 309.8 0.133
PET [37] ICCV’23 74.4 328.5 -
APGCC [38] ECCV’24 71.7 284.4 0.215
Gramformer [39] AAAI’24 72.5 316.4 0.160
mPrompt [18] CVPR’24 62.1 293.5 -
CLIP-EBC [15] ICME’25 61.3 278.4 0.148
EBC-ZIP-B (ours) - 60.1 299.0 0.104
TABLE V: Comparison of our method EBC-ZIP-B with the latest approaches on the test split of NWPU-Crowd. EBC-ZIP achieves both the lowest MAE and NAE on this large-scale benchmark.
IV-B EBC-ZIP-B Compared with State-of-the-Art
We first compare our proposed model, EBC-ZIP-B (the backbone choice and computational complexity can be found in Table I), with recent state-of-the-art crowd counting methods on four benchmark datasets: ShanghaiTech A & B, UCF-QNRF, and the validation split of NWPU-Crowd. The results are summarized in Table IV. Our method consistently achieves the best performance across all datasets. In particular, on ShanghaiTech A, our model achieves the lowest MAE (47.81) and RMSE (75.04), outperforming the previous best result (48.8 MAE, 76.7 RMSE by APGCC [38]). On ShanghaiTech B, we also achieve the best performance with a MAE of 5.51 and RMSE of 8.63, significantly surpassing strong baselines such as CrowdHat [36] and STEERER [13]. For the challenging UCF-QNRF dataset, our method sets a new state-of-the-art with 69.46 MAE and 121.88 RMSE. On NWPU-Crowd (val), we obtain a substantial improvement over prior methods, achieving 28.26 MAE and 64.84 RMSE, significantly outperforming the closest competitor, CLIP-EBC [15], which reports 36.6 MAE and 81.7 RMSE.
We further evaluate our method on the NWPU-Crowd test set, with results summarized in Table V. Our EBC-ZIP-B achieves the lowest MAE (60.1) and lowest NAE (0.104) among all methods, indicating superior overall accuracy and robustness. Although the RMSE (299.0) is not the lowest, it is comparable to other top-performing methods. Notably, RMSE tends to amplify the influence of outlier errors. Given our observation of annotation noise, such as duplicate or out-of-bound head labels in the training and validation splits, RMSE may not always reflect true performance. In contrast, MAE and NAE offer more stable and representative evaluations, making our results more reliable in practice. These results clearly demonstrate the effectiveness and generalizability of our method across various crowd densities and scenes.
ShanghaiTech A ShanghaiTech B UCF-QNRF
Method Size (M) FLOPS (G) MAE RMSE MAE RMSE MAE RMSE
LSANet [25] 0.20 14.40 66.1 110.2 8.6 13.9 112.3 186.9
EBC-ZIP-P (ours) 0.81 6.46 71.18 109.6 8.23 12.62 96.29 161.82
EBC-ZIP-N (ours) 3.36 24.73 60.21 95.61 7.74 12.14 86.46 147.64
MobileCount [16] 3.40 16.49 89.4 146.0 9.0 15.4 131.1 222.6
LMSFFNet-XS [26] 4.58 39.28 85.8 139.9 9.2 15.1 112.8 201.6
MobleCount×$\times 1.25$ [16] 5.47 34.07 82.9 137.9 8.2 13.2 124.5 207.6
EBC-ZIP-T (ours) 10.53 61.39 59.07 90.55 6.67 9.90 76.02 129.40
MobleCount×$\times 2$ [16] 13.39 63.03 81.4 133.3 8.1 12.7 117.9 207.5
EBC-ZIP-S (ours) 33.60 242.43 57.20 93.07 5.88 9.37 73.32 125.09
TABLE VI: Comparison of our lightweight models with state-of-the-art lightweight approaches on ShanghaiTech A & B, and UCF-QNRF. Models are grouped by parameter size: $<1$ M, $1-10$ M, $10-20$ M, $\geq 20$ M. FLOPs are measured on inputs of size $1920\times 1080$. The best result in each group is highlighted, the overall best result is shown in bold, and the second-best is underlined.
Suffix MAE M1Pro CPU M1Pro GPU 5900X 3090
-P 8.23 9.99 118.63 122.22 1,012.01
-N 7.74 5.28 63.17 50.67 548.14
-T 6.67 1.87 27.77 18.37 249.73
-S 5.88 0.48 6.80 4.60 50.91
TABLE VII: Inference speed (frames per second, FPS) of the lightweight variants on ShanghaiTech B ($1024\times 768$ resolution). Results are reported on four hardware platforms.
IV-C Comparison with Lightweight Models
To investigate the performance-efficiency trade-off, we compare our proposed lightweight models (EBC-ZIP-P, EBC-ZIP-N, EBC-ZIP-T, and EBC-ZIP-S) with a range of representative compact crowd counting methods. The backbones of these variants can be found in Table I. We divide models into four groups based on parameter size: (1) ultra-lightweight models ($<$1M), (2) lightweight models (1–10M), (3) mid-size models (10–20M), and (4) larger lightweight models ($\geq$20M). Table VI presents the details.
Group I: Ultra-lightweight ($<$1M).
In this group, EBC-ZIP-P achieves the best performance across all three benchmarks, with notably lower MAE on ShanghaiTech B (8.23) and UCF-QNRF (96.29) than LSANet [25], despite having over 50% fewer FLOPs. This demonstrates our model’s superior design in extremely constrained settings.
Group II: Lightweight (1–10M).
EBC-ZIP-N significantly outperforms other methods such as MobileCount [16] and LMSFFNet-XS [26] across all datasets. It achieves the lowest MAE in this group on ShanghaiTech A (60.21), ShanghaiTech B (7.74), and UCF-QNRF (86.46).
Group III: Mid-size (10–20M).
EBC-ZIP-T leads this group with the best results on all benchmarks. It achieves group-best scores of 59.07 MAE on ShanghaiTech A, 6.67 on ShanghaiTech B, and 76.02 on UCF-QNRF, significantly outperforming MobileCount×$\times$2 [16] with fewer parameters and FLOPs.
Group IV: Larger lightweight ($\geq$20M).
EBC-ZIP-S sets a new benchmark in this group and across all lightweight models, achieving 57.20 MAE on ShanghaiTech A, 5.88 on ShanghaiTech B, and 73.32 on UCF-QNRF. These scores represent the best results among all compared lightweight models, and they also approach or surpass those of regular-sized models presented in Table IV.
Across all groups, our EBC-ZIP variants consistently outperform existing methods with comparable or fewer parameters and FLOPs. These results validate the effectiveness and scalability of our EBC-ZIP framework, making it suitable for real-world deployments under various computational constraints.
To assess the practical deployment potential of our lightweight models, we measure the average inference speed of each EBC-ZIP variant across four representative hardware platforms: Apple M1 Pro CPU (6 performance + 2 efficiency cores), Apple M1 Pro 14-core GPU, AMD Ryzen 9 5900X CPU (12-core), and NVIDIA RTX 3090 GPU. All experiments are conducted on the ShanghaiTech B dataset using full-precision inference at an input resolution of $1024\times 768$, which reflects real-world surveillance camera scenarios. The PyTorch version is 2.7.1. Table VII reports the average frames per second (FPS) for each model. Our smallest variant, EBC-ZIP-P, achieves the highest throughput on all platforms, reaching 122.22 FPS on AMD 5900X and over 1,000 FPS on RTX 3090, making it ideal for resource-constrained or real-time applications. On the other end, EBC-ZIP-S delivers the best accuracy (MAE = 5.88) but at a higher computational cost, with CPU-side inference limited to under 5 FPS, making it suitable for accuracy-critical but latency-tolerant use cases. EBC-ZIP-N and EBC-ZIP-T offer compelling accuracy-efficiency trade-offs. For example, EBC-ZIP-T achieves 18.37 FPS on AMD 5900X and 249.73 FPS on RTX 3090, while maintaining a competitive MAE of 6.67. These results indicate that mid-sized variants can enable near real-time performance even on high-end CPUs without requiring dedicated GPUs. Overall, our EBC-ZIP family provides a scalable solution space, allowing deployment under varying hardware and latency constraints without compromising on count accuracy.
Backbone: VGG19 Training Resolution
448 672 896
Block Size 16 78.52 75.06 77.92
32 75.73 73.02 75.69
TABLE VIII: Ablation study of training resolution and block size on UCF-QNRF with the VGG19-based backbone. Performances are measured by MAE.
IV-D Influence of Training Resolution and Block Size
We empirically justify our choice of a training resolution of 672 pixels and a block size of 32 pixels on UCF-QNRF through a two-factor ablation study with a VGG-19 backbone. The resulting MAE values are reported in Table VIII. A resolution of 672 pixels consistently outperforms 448 pixels and 896 pixels for both block sizes, and a larger block size of 32 pixels yields lower errors than 16 pixels at every resolution. The combination (672 pixels, 32 pixels) attains the overall best MAE of 73.02, and we therefore adopt this setting for all subsequent UCF-QNRF experiments.
IV-E Visualization
(a) Input Image.
(b) GT Density Map
(c) Structural Zero Map
(d) Predicted Density Map
Figure 6: Qualitative example of EBC-ZIP-B on a ShanghaiTech B test image. The structural-zero map (red = high probability of background/body, blue = head-center region) automatically suppresses background and non-head body areas while preserving only the small central region around each head annotation, resulting in a more accurate density estimate.
Figure 6 presents a moderate-density street-view scene from the test set of ShanghaiTech B together with the corresponding EBC-ZIP outputs. The structural-zero map in 6(c) predicts nearly all non-head pixels (including walls, torso regions, and pavement) as zeros, while preserving only the compact blue blobs centered on annotated heads. This confirms that the $\boldsymbol{\pi}$-head behaves as an implicit, fine-grained foreground segmenter. Leveraging this mask, the predicted density in Fig.6(d) cleanly localizes every head and reproduces the ground-truth count (31) with virtually no background false positives. These results demonstrate that explicit modeling of structural zeros directs the network’s attention to true head centers and improves counting accuracy.
V Conclusion
We presented EBC-ZIP, a crowd-counting framework that improves the Enhanced Blockwise Classification (EBC) framework by integrating zero-inflated Poisson (ZIP) regression. By learning a structural-zero branch to suppress background and non-head body regions, and a Poisson-rate branch to model true head centers, the network tackles the extreme sparsity of density maps without any explicit segmentation supervision. Comprehensive experiments show that the full model, EBC-ZIP-B, establishes new state-of-the-art MAE on ShanghaiTech A & B, UCF-QNRF, and NWPU-Crowd, and lowers NAE on the NWPU test set by 21.8% relative to the best previous method. We further introduce a family of lightweight variants (from Pico to Small) that retain these accuracy gains under tight computational budgets and consistently outperform existing models of comparable size on all four benchmarks, highlighting both the effectiveness and scalability of our EBC-ZIP framework. Future work will extend EBC-ZIP to video-based crowd counting and explore additional temporal cues.