### v2

# ZIP: Scalable Crowd Counting via Zero‑Inflated Poisson Modeling

Abstract
Most crowd counting methods directly regress blockwise density maps using Mean Squared Error (MSE) losses. This practice has two key limitations: 1) it fails to account for the extreme spatial sparsity of annotations—over 95% of $8\times 8$ blocks are empty across standard benchmarks, so supervision signals in informative regions are diluted by the predominant zeros; 2) MSE corresponds to a Gaussian error model that poorly matches discrete, non‑negative count data. To address these issues, we introduce ZIP, a scalable crowd counting framework that models blockwise counts with a Zero‑Inflated Poisson likelihood: a zero‑inflation term learns the probability a block is structurally empty (handling excess zeros), while the Poisson component captures expected counts when people are present (respecting discreteness). We provide a generalization analysis showing a tighter risk bound for ZIP than MSE‑based losses and DMCount provided that the training resolution is moderately large. To assess the scalability of ZIP, we instantiate it on backbones spanning over 100×100$\times$100 × in parameters/compute. Experiments on ShanghaiTech A & B, UCF-QNRF, and NWPU-Crowd demonstrate that ZIP consistently surpasses state‑of‑the‑art methods across all model scales.
Introduction
Crowd counting aims to estimate the number of people present in an image or a video frame. It underpins a broad range of real-world applications, including public safety and crowd management (Valencia et al. 2021; Saleh, Suandi, and Ibrahim 2015), urban planning (Waclawek et al. 2024), and intelligent transportation systems (McCarthy et al. 2025).
Most modern crowd counting methods (Ma et al. 2019; Cheng et al. 2022; Han et al. 2023; Ranasinghe et al. 2024) construct blockwise training targets by first convolving point annotations with Gaussian kernels to create a pixel‑level density map, and then summing density values within each $B\times B$ block (e.g., $8\times 8$) to obtain blockwise density maps. Models are trained to regress these blockwise targets (usually with Mean Squared Error (MSE)) in lieu of raw pixel supervision to mitigate extreme pixel‑level sparsity.
In these methods, however, the aggregated blockwise density maps remain highly sparse: across ShanghaiTech A & B (Zhang et al. 2016), UCF‑QNRF (Idrees et al. 2018), and NWPU‑Crowd (Wang et al. 2020c), over 95% of $8\times 8$ blocks contain zero people (empirical measurement). This zero dominance skews squared‑error losses, dilutes supervision from populated regions, and biases models toward under‑counting. These methods also suffer label noise introduced by Gaussian smoothing. Point annotations are spatially uncertain: a human click marks an approximate head center, and nearby pixels could be equally valid. Spreading each point with a local kernel is intended to make supervision tolerant to such subjective variation. However, since the true head extents are unknown, mismatched kernel bandwidths yield over‑ or under‑counts after block aggregation and thus inject noise into learning targets.
(a) Input Image.
(b) GT Density Map.
(c) Structural Zero Map.
(d) Predicted Density Map.
Figure 1: Visualization of ZIP on a ShanghaiTech B test image. Color scale (all panels): red = higher value, blue = lower value (per-panel normalization). In Fig. 1(c), red regions have high probability of being structurally empty (background, torso, peripheral head parts), blue indicates low structural-zero probability (candidate head centers). In Fig. 1(d), red regions correspond to high expected local crowd density near head centers; blue denotes near-zero density.
Prior work has tried to mitigate sparsity by first segmenting foreground crowd regions and then counting only within those areas (Rong and Li 2021; Modolo et al. 2021; Guo et al. 2024). Such a practice introduces additional multi-task loss balancing complexity. Also, since crowd counting datasets are not annotated with segmentation masks, these approaches usually derive pseudo masks from Gaussian-smoothed density maps, inheriting the same head-size uncertainty and label noise discussed above.
Figure 2: Illustration of structural and sampling zeros (best viewed in color). The image (synthesized for illustration purposes) is overlaid with a grid where each cell represents a spatial block. The red dot marks a ground-truth head annotation, and the block containing it is labeled with the red number “1”. Yellow zeros, which indicate sampling zeros, are assigned to blocks that correspond to the head-center region but receive zero count due to the point-based annotation protocol. Green zeros denote structural zeros, corresponding to background regions, non-head body parts or outer regions of the head that are not associated with any annotations.
Additionally, MSE-based loss functions present a modeling mismatch. Blockwise targets represent discrete, non‑negative, and zero-heavy count data, whereas MSE implicitly assumes Gaussian residuals on continuous values. A Poisson model captures both the discreteness and non-negativity of counts, yet a standard Poisson struggles with the overwhelming presence of zeros. These gaps motivate a formulation that explicitly separates structural emptiness from stochastic counts, leading to our Zero-Inflated Poisson (ZIP) framework.
Our ZIP framework bypasses Gaussian smoothing and model per-block counts $y\in\mathbb{N}$ via a zero‑inflated Poisson:
$P(y=k|\pi,\lambda)=\begin{cases}\pi+(1-\pi)e^{-\lambda},\quad&k=0\\ (1-\pi)\frac{e^{-\lambda}\lambda^{k}}{k!},\quad&k>0\end{cases}$ (1)
with $\pi$, $\lambda$ predicted per block. In this way, we hypothesize that zeros in blockwise density maps arise from two distinct mechanisms (See Fig. 2 for illustration):
  • Structural zeros (modeled by $\pi$), which are blocks that are deterministically zero due to their semantic irrelevance. These include background (sky, buildings, etc.), body parts (limbs, chest, etc.) and peripheral head regions that do not correspond to head centers, and they comprise the majority of zero blocks.
  • Sampling zeros (modeled by $(1-\pi)e^{-\lambda}$): each head is annotated by a single point that we assign to exactly one supervision block. Neighboring blocks that also correspond to the head center therefore receive zero count. Modeling these annotation effects in the Poisson component absorbs the positional ambiguity that Gaussian smoothing was meant to solve, removing kernel bandwidth tuning.
Because structural and sampling zeros are not annotated separately, our ZIP model learns to distinguish these two types of zeros implicitly by optimizing the negative log-likelihood of the ZIP distribution, with $\lambda$ additionally supervised by a cross entropy loss to stabilize training. Empirically, learned zero‑inflation correlates with non-head-center areas. As illustrated in Fig. 1, the structural-zero map in 1(c) predicts nearly all non-head pixels as zeros, while preserving only the compact blue blobs centered on heads. Compared with segmentation-based methods, our ZIP-based method offers several notable advantages: 1) it does not rely on segmentation masks, thereby avoiding noise introduced by Gaussian smoothing; 2) it captures structural sparsity without assuming the presence of clear object boundaries, making it better suited for highly occluded and densely crowded scenes; 3) it eliminates the need to balance counting and segmentation, simplifying the training objective.
Figure 3: Scalability of the proposed ZIP framework. The $x$-axis represents the number of model parameters (in millions), and the $y$-axis indicates the Mean Absolute Error (MAE; lower is better) on the validation set of the ShanghaiTech B dataset (Zhang et al. 2016). Each circle represents a specific model, where the radius corresponds to the number of floating point operations (FLOPs) required to process a single $1920\times 1080$ image during inference. ZIP models (in blue) exhibit a favorable trade-off between model size and accuracy, demonstrating better scalability compared to other lightweight (yellow) and heavyweight (green) models.
Through theoretical risk bound analysis (See Theorem 1) and comprehensive empirical evaluation, we show that ZIP consistently outperforms Gaussian-smoothed MSE-based losses and DMCount (Wang et al. 2020a). To further assess the scalability of ZIP, we evaluate its performance with backbones of varying computational complexities, including MobileNet, MobileCLIP , and CLIP-ConvNeXt. As illustrated in Fig. 3, our ZIP framework consistently outperforms existing models in both accuracy (measured by MAE) and computational efficiency (measured by FLOPs) across different parameter scales, demonstrating its superior scalability. To summarize, our contributions are as follows:
  • We propose a novel crowd counting framework ZIP based on Zero-Inflated Poisson (ZIP) regression, which explicitly models the extreme sparsity of ground-truth count distributions by disentangling structural zeros from sampling uncertainty.
  • We conduct a theoretical analysis and experiments on four crowd counting benchmarks to validate the effectiveness of our ZIP framework. Results show that our base model, ZIP-Base, consistently outperforms existing state-of-the-art methods across four different datasets.
  • We present a systematic study of framework-level scalability in crowd counting. Our ZIP framework generalizes well across a wide range of model sizes and architectures, from lightweight convolutional neural networks (CNNs) to vision-language models, achieving state-of-the-art performance under varying computational constraints.
Related Work
To address the extreme sparsity of ground-truth density maps, many existing methods (Zhang et al. 2016; Liu, Salzmann, and Fua 2019; Ma et al. 2019; Wang et al. 2020b, 2021; Han et al. 2023; Guo et al. 2024) adopt Gaussian smoothing. These approaches typically preprocess the ground-truth density map by convolving it with Gaussian kernels, either of fixed or adaptive size. In principle, the kernel size should reflect the actual head size in the image; however, such information is typically unavailable in most crowd counting datasets. As a result, Gaussian smoothing inevitably introduces noise into the ground-truth density maps, which can degrade the model’s generalization ability (Wang et al. 2020a). There have also been some efforts (Rong and Li 2021; Modolo et al. 2021; Guo et al. 2024) to address the sparsity by training a segmentation model which aims to separate regions containing people from the background. However, most crowd counting datasets do not provide segmentation masks, they have to rely on Gaussian-smoothed ground-truth density maps to generate pseudo ground-truth segmentation masks. This again introduces noise into the pseudo segmentation masks. Guo et al. exploited the box annotations provided by NWPU-Crowd (Wang et al. 2020c) to handle this problem, but compared with point annotations, box annotations are much more expensive to acquire. Besides, these methods introduce an extra task to counting, and how to balance the loss terms in such a multi-task setting poses a new challenge.
Figure 4: Overview of the proposed ZIP framework. Given an input image $\boldsymbol{X}$ of spatial size $H\times W$, a backbone network extracts feature maps, which are shared by two parallel branches. In the Poisson branch (top), the features are passed to a $\boldsymbol{\lambda}$-head with a softmax activation, producing the blockwise count distribution $\boldsymbol{P}^{*}_{\boldsymbol{\lambda}}$ over $(n-1)$ positive bins. The Poisson rate per block is then computed as a weighted average of bin centers, yielding the final $\boldsymbol{\lambda}$ map. In the Zero Inflation branch (bottom), the same features are fed into a $\boldsymbol{\pi}$-head with a sigmoid activation to estimate the structural zero probability map $\boldsymbol{\pi}$ of spatial size $h\times w$. The predicted density map $\boldsymbol{Y}^{*}$ is defined as the expected value of the zero-inflated Poisson distribution: $\boldsymbol{Y}^{*}=(1-\boldsymbol{\pi})\otimes\boldsymbol{\lambda}$, where $\otimes$ denotes elementwise multiplication.
On the other hand, Most density-based crowd counting methods typically frame the task as a regression problem, where the model is trained to minimize the blockwise mean squared error (MSE) between the predicted and ground-truth blockwise density maps (Zhang et al. 2016; Liu, Salzmann, and Fua 2019; Liu et al. 2019; Ma et al. 2019; Wang et al. 2020b; Han et al. 2023; Guo et al. 2024). This loss function implicitly assumes that each blockwise count follows a Gaussian distribution centered at the predicted value. This formulation fails to capture the discreteness and non-negativity of count data. DMCount (Wang et al. 2020a) addresses a different issue by reformulating density estimation as an Earth Mover’s Distance (EMD) problem. By minimizing the Wasserstein distance between predicted and ground-truth density maps, DMCount explicitly models the global transportation cost needed to match the predicted mass to the annotated points. Similar to MSE-based methods, it still operates entirely in a continuous density space, treating the density maps as divisible mass rather than integer-valued counts. Also, the exact computation of EMD scales poorly with the number of bins ($\mathcal{O}(n^{3})$).
Method
In this work, we propose to incorporate Zero-Inflated Poisson (ZIP) regression to address the extreme sparsity of ground-truth density maps. Fig. 4 illustrates an overview of the proposed ZIP framework.
Zero-Inflated Poisson Regression
Since training crowd counting models directly via regression can suffer from unstable gradient updated, we follow the Enhanced Block Classification (EBC) framework (Ma, Sanchez, and Guha 2024) to first quantize blockwise counts into integer-valued bins $\{\mathcal{B}_{k}\}_{k=1}^{n}$ with bin centers $\{b_{k}\}_{k=1}^{n}$.
(a) MAE
(b) RMSE
(c) NAE (%)
Figure 5: Performance of the VGG19-based structure on ShanghaiTech B under varying values of $\omega$ in the loss Eq. (11). Results indicate that $\omega=0.75$ yields the best MAE (6.34) and RMSE (9.79), while $\omega=1.00$ gives the lowest NAE (5.00%).
Given an input image $\boldsymbol{X}\in\mathbb{R}^{3\times H\times W}$, we first compute the feature map $\boldsymbol{f}$ using a shared backbone:
$\boldsymbol{f}=\mathcal{F}(\boldsymbol{X})\in\mathbb{R}^{C\times h\times w}$ (2)
where $h=H//r$, $w=W//r$ denote the spatial size of the output.
Then, the feature map $\boldsymbol{f}$ is passed through a $\boldsymbol{\lambda}$-head $\mathcal{H}_{\boldsymbol{\lambda}}$ (implemented as a $1\times 1$ convolution), followed by a softmax activation to generate a probability distribution over the $n{-}1$ positive count bins:
$\boldsymbol{P}^{*}_{\boldsymbol{\lambda}}=\mathrm{Softmax}\left(\mathcal{H}_{\boldsymbol{\lambda}}(\boldsymbol{f})\right)\in\mathbb{R}^{(n-1)\times h\times w}.$ (3)
Note that we exclude the $\{0\}$ bin here, since the Poisson distribution assumes a strictly positive rate $\lambda>0$. We then compute the estimated Poisson rate map $\boldsymbol{\lambda}\in\mathbb{R}_{+}^{1\times h\times w}$ as the expected value of the bin centers:
$\boldsymbol{\lambda}_{i,j}=\sum_{k=2}^{n}{\boldsymbol{P}^{*}_{\boldsymbol{\lambda}}}_{k,i,j}\cdot{b}_{k},$ (4)
where ${b}_{k}$ for $k=2,\cdots,n$ denotes the center of each positive bin.
The learning of $\boldsymbol{P}^{*}_{\boldsymbol{\lambda}}$ is supervised by a cross-entropy loss over the positive regions of the ground-truth density map. Specifically, we define the set of positive-valued blocks as $\boldsymbol{Y}_{+}\coloneqq\boldsymbol{Y}[\boldsymbol{Y}>0]$, which arise from the Poisson component and are thus only related to $\boldsymbol{\lambda}$. Their corresponding probabilistic predictions $\boldsymbol{P}^{*}_{\boldsymbol{\lambda}+}$ can be obtained via $\boldsymbol{P}^{*}_{\boldsymbol{\lambda}+}=\boldsymbol{P}^{*}_{\boldsymbol{\lambda}}[:,\boldsymbol{Y}>0]$ The cross-entropy loss is calculated based on $\boldsymbol{P}^{*}_{\boldsymbol{\lambda}+}$ and $\boldsymbol{Y}_{+}$:
$\mathcal{L}_{\mathrm{CE}}=\mathrm{CrossEntropy}(\boldsymbol{P}^{*}_{\boldsymbol{\lambda}+},\boldsymbol{P}_{+}),$ (5)
where $\boldsymbol{P}_{+}$ is the one-hot encoded ground-truth probability map of $\boldsymbol{Y}_{+}$.
In parallel, we introduce a $\boldsymbol{\pi}$-head, $\mathcal{H}_{\boldsymbol{\pi}}$, also implemented as a $1\times 1$ convolution, followed by a sigmoid activation to produce the structural zero probability map:
$\boldsymbol{\pi}=\sigma\left(\mathcal{H}_{\boldsymbol{\pi}}(\boldsymbol{f})\right)\in\mathbb{R}^{1\times h\times w}$ (6)
The negative log-likelihood (NLL) loss of the zero-inflated Poisson distribution is given by:
$\mathcal{L}_{\mathrm{NLL}}=-\sum_{i=1}^{h}\sum_{j=1}^{w}\log P(\boldsymbol{Y}_{i,j}\mid\boldsymbol{\pi}_{i,j},\boldsymbol{\lambda}_{i,j})$ (7)
where $P(\boldsymbol{Y}_{i,j}\mid\boldsymbol{\pi}_{i,j},\boldsymbol{\lambda}_{i,j})$ is the p.m.f. of the ZIP distribution given by Eq. (1), and $\boldsymbol{\lambda}$ and $\boldsymbol{\pi}$ are given by (4) and (6), respectively.
It is important to note that we do not supervise the $\boldsymbol{\pi}$ and $\boldsymbol{\lambda}$ branches separately. Since structural and sampling zeros are not explicitly annotated in the ground-truth, we treat them as latent factors and optimize both branches jointly via the ZIP negative log-likelihood loss Eq. (7). This formulation allows the model to implicitly learn to disentangle the two types of zeros based on the statistical patterns in the data, without requiring additional annotations.
We use the expectation map $\boldsymbol{Y}^{*}=\mathbb{E}[\boldsymbol{V}|\boldsymbol{\pi},\boldsymbol{\lambda}]$ as the predicted density map, given by
$\boldsymbol{Y}^{*}_{i,j}=\mathbb{E}[\boldsymbol{V}_{i,j}\mid\boldsymbol{\pi}_{i,j},\boldsymbol{\lambda}_{i,j}]=(1-\boldsymbol{\pi})\otimes\boldsymbol{\lambda},$ (8)
where $\otimes$ represents element-wise multiplication.
To ensure accurate crowd estimates at the image level, we incorporate a count loss that penalizes discrepancies between the predicted and ground-truth total counts. The predicted total count $c^{*}$ is obtained by summing over all elements of the predicted density map:
$c^{*}=\sum_{i=1}^{h}\sum_{j=1}^{w}\boldsymbol{Y}^{*}_{i,j}.$ (9)
We define the count loss as the mean absolute error (MAE) between the predicted and ground-truth total counts:
$\mathcal{L}_{\mathrm{count}}=\left|c^{*}-c\right|=\left|\sum_{i=1}^{h}\sum_{j=1}^{w}\left(\boldsymbol{Y}^{*}_{i,j}-\boldsymbol{Y}_{i,j}\right)\right|$ (10)
Finally, the model is trained using a weighted sum of three loss terms:
$\mathcal{L}_{\mathrm{total}}=\omega\mathcal{L}_{\mathrm{CE}}+\mathcal{L}_{\mathrm{NLL}}+\mathcal{L}_{\mathrm{count}},$ (11)
where $\omega$ is a scalar weighting coefficients, and $\mathcal{L}_{\mathrm{CE}}$, $\mathcal{L}_{\mathrm{NLL}}$, and $\mathcal{L}_{\mathrm{count}}$ are defined in Eq. (5), (7), and (10), respectively.
ZIP Has a Tighter Risk Bound
Since Wang et al. established that DMCount enjoys a tighter risk bound than Gaussian-smoothed MSE-based losses, it suffices for us to show that the ZIP NLL loss defined in Eq. (7) admits an even tighter risk bound than DMCount. We formalize this result in the following theorem.
Theorem 1.
Assume the per-block counts follow a zero-inflated Poisson law with $\boldsymbol{\pi}>0$. Let $\mathcal{D}$ be the unknown ground-truth data distribution and $\mathcal{S}$ be the i.i.d. $K$ samples. Denote the hypothesis class of scalar regressors by $\mathcal{H}$ and the family of vector blockwise density predictors by =×$\mathcal{F}=\mathcal{H}^{\times hw}$. Let $f_{\mathcal{S}}^{\mathrm{NLL}}$ denotes an empirical risk minimizer of $\mathcal{L}_{\mathrm{NLL}}$ on $\mathcal{S}$, and $f_{\mathcal{D}}$ is the population minimizer (“Bayes optimal predictor”) for the same loss. Then, for any $\delta\in(0,1)$, with probability at least $1-\delta$ over the choice of the $K$ i.i.d. training examples $\mathcal{S}$, every $f\in\mathcal{F}$ satisfies
$\mathcal{R}_{\mathcal{D}}(f_{\mathcal{S}}^{\mathrm{NLL}},\mathcal{L}_{\mathrm{NLL}})-\mathcal{R}_{\mathcal{D}}(f_{\mathcal{D}},\mathcal{L}_{\mathrm{NLL}})$
$\leq 2\cdot hw\cdot C\cdot R_{\mathcal{S}}(\mathcal{H})+\mathcal{O}(\sqrt{1/K}),$ (12)
where $hw$ is the total number of blocks, $C$ is a dataset-specific constant, and $R_{\mathcal{S}}(\mathcal{H})$ is the empirical Rademacher complexity of the hypothesis class $\mathcal{H}$ on the sample set $\mathcal{S}$.
By comparison, generalization gap of the OT loss in DMCount (Wang et al. 2020a) is bounded by
$\mathcal{R}_{\mathcal{D}}(f_{\mathcal{S}}^{\mathrm{OT}},\mathcal{L}_{\mathrm{OT}})-\mathcal{R}_{\mathcal{D}}(f_{\mathcal{D}},\mathcal{L}_{\mathrm{OT}})$
$\leq$ $4\cdot(hw)^{2}\cdot C_{\infty}\cdot R_{\mathcal{S}}(\mathcal{H})+\mathcal{O}(\sqrt{1/K}).$ (13)
Hence, when $h$ or $w$ grows, the upper bound of ZIP (Eq. (12)) becomes tighter than that of OT (Eq. (13)). The proof of Theorem 1 can be found in the supplementary material.
Experiments
Variants. To evaluate the scalability and adaptability of our proposed framework, we construct five model variants of ZIP with different backbones, including MobileNetV4 (Qin et al. 2024), MobileCLIP (Vasu et al. 2024) and OpenCLIP’s ConvNeXt-Base (Liu et al. 2022; Cherti et al. 2023) These variants are suitable for different application scenarios, ranging from mobile-friendly to high-performance settings. The specifications are summarized in Table 1. All variants share the same head design and differ only in backbone complexity. This modularity allows ZIP to scale across computational budgets.
Setup. Our training configuration largely follows. We first preprocess all datasets so that the minimum edge length is no less than 448. We limit the maximum edge length of UCF-QNRF to be 2048 and that of NWPU-Crowd to be 3072. We apply a combination of data augmentation techniques, including random resized cropping, horizontal flipping, color jittering, and salt-and-pepper noise injection. For ShanghaiTech A and B, which consist of lower-resolution images, the crop size is set to $448\times 448$. For the higher-resolution datasets UCF-QNRF and NWPU-Crowd, we use a larger crop size of $672\times 672$. All models are optimized using the Adam optimizer with a weight decay of $41e-$ 4. The learning rate is linearly warmed up from $51e-$ 5 to $41e-$ 4 over the first 25 epochs, followed by cosine annealing with parameters $T_{0}=5$ and $T_{\mathrm{mult}}=2$.
Evaluation metrics. Following standard practice, we evaluate each model variant on all datasets using three commonly adopted metrics: Mean Absolute Error (MAE), Root Mean Squared Error (RMSE), and Normalized Absolute Error (NAE). The NAE metric ensures fair evaluation across both sparse and dense scenes by normalizing the absolute error with respect to the ground-truth count. These metrics are defined as follows:
$\mathrm{MAE}=\frac{1}{M}\sum_{m=1}^{M}\left|c_{m}-c_{m}^{*}\right|,$ (14)
$\mathrm{RMSE}=\sqrt{\frac{1}{M}\sum_{m=1}^{M}\left|c_{m}-c_{m}^{*}\right|^{2}},$ (15)
and
$\mathrm{NAE}=\frac{1}{\|\mathcal{M}_{+}\|}\sum_{m\in\mathcal{M}_{+}}\frac{\left|c_{m}-c_{m}^{*}\right|}{c_{m}},$ (16)
where $M$ denotes the total number of test images, $c_{m}$ is the ground-truth count for image $m$, and $c_{m}^{*}$ is the predicted count, computed via Eq. (9). The set $\mathcal{M}+\subseteq{1,2,\dots,M}$ includes only the indices for which $c_{m}>0$, and $\|\mathcal{M}_{+}\|$ denotes its cardinality.
Suffix Backbone Size (M) FLOPS (G)
-P (Pico) MobileNetV4-Small-0.5×0.5$\times$0.5 × 0.81 6.46
-N (Nano) MobileNetV4-Small 3.36 24.73
-T (Tiny) MobileNetV4-Medium 10.64 61.62
-S (Small) MobileCLIP-S1 33.72 242.68
-B (Base) CLIP-ConvNeXt-Base 105.60 800.99
Table 1: ZIP variants with corresponding backbones, parameter sizes, and computational complexity measured on $1920\times 1080$ resolution.
Component MAE RMSE NAE
MAE 6.69 10.45 5.52%
DMCount (Wang et al. 2020a) 6.62 9.83 5.30%
Poisson NLL 6.57 9.86 5.47%
$\mathcal{L}_{\mathrm{NLL}}$ (ours) 6.36 10.32 5.00%
Table 2: Comparison of $\mathcal{L}_{\mathrm{NLL}}$ (Eq. (7)) with alternative choices in the total loss function (Eq. (11)) on ShanghaiTech B, while $\mathcal{L}_{\mathrm{CE}}$ and $\mathcal{L}_{\mathrm{count}}$ remain unchanged. Results show that our ZIP NLL achieves the lowest MAE and NAE.
Method ShanghaiTech A ShanghaiTech B UCF-QNRF NWPU (Val)
MAE RMSE MAE RMSE MAE RMSE MAE RMSE
ChfL (Shu et al. 2022) 57.5 94.3 6.9 11.0 80.3 137.6 76.8 343.0
MAN (Lin et al. 2022) 56.8 90.3 - - 77.3 131.5 76.5 323.0
CLTR (Liang, Xu, and Bai 2022) 56.9 95.2 6.5 10.6 85.8 141.3 61.9 246.3
CrowdHat (Wu and Yang 2023) 51.2 81.9 5.7 9.4 75.1 126.7 - -
STEERER (Han et al. 2023) 54.5 86.9 5.8 8.5 74.3 128.3 54.3 238.3
PET (Liu et al. 2023) 49.3 78.8 6.2 9.7 79.5 144.3 58.5 238.0
APGCC (Chen et al. 2024) 48.8 76.7 5.6 8.7 80.1 133.6 - -
Gramformer (Lin et al. 2024) 54.7 87.1 - - 76.7 129.5 - -
mPrompt (Guo et al. 2024) 52.5 88.9 5.8 9.6 72.2 133.1 50.2 219.0
CLIP-EBC (Ma, Sanchez, and Guha 2024) 52.5 83.2 6.0 10.1 80.3 136.6 36.6 81.7
ZIP-B (ours) 47.81 75.04 5.51 8.63 69.46 121.88 28.26 64.84
Table 3: Comparison of our model ZIP-B with state-of-the-art approaches on ShanghaiTech A & B, UCF-QNRF, and the NWPU-Crowd validation split. ZIP-B achieves the best performance across all these four datasets under both MAE and RMSE, surpassing all strong baselines.
Effectiveness of ZIP
In this subsection, we empirically validate the effectiveness of our proposed ZIP framework by comparing it against several strong baselines, including BL (Ma et al. 2019), DMCount (Wang et al. 2020a), and EBC (Ma, Sanchez, and Guha 2024). For a fair comparison, we adopt a consistent encoder-decoder architecture based on VGG19 (Simonyan and Zisserman 2015). Before proceeding to the comparison, we first determine the optimal value of the weighting parameter $\omega$ in Eq.(11), which balances the cross-entropy term and other two terms. To this end, we evaluate the VGG19-based model under various settings of $\omega$ on the ShanghaiTech B dataset and report the results in Fig. 5. As illustrated, the model achieves the best MAE and RMSE when $\omega=0.75$, and the lowest NAE when $\omega=1.00$. In our subsequent comparisons, we fix $\omega=1.00$ to prioritize NAE, due to its robustness to varying crowd densities.
To validate the effectiveness of the proposed ZIP NLL loss, we conduct an ablation study where we replace the ZIP NLL component in the total loss function (11) with several commonly used alternatives, including MAE, DMCount (Wang et al. 2020a), and Poisson NLL. Other components in the loss function, such as the cross-entropy loss and the count loss, are kept unchanged. All models are trained using the same VGG19-based architecture under identical settings. The evaluation results on ShanghaiTech B are presented in Table 2. As shown, our ZIP NLL achieves the lowest MAE (6.36) and NAE (5.00%), and performs competitively in RMSE (10.32), demonstrating its effectiveness in enhancing counting accuracy while maintaining robust distribution modeling.
Method MAE RMSE NAE
ChfL (Shu et al. 2022) 76.8 343.0 0.171
MAN (Lin et al. 2022) 76.5 323.0 0.170
CLTR (Liang, Xu, and Bai 2022) 74.4 333.8 0.165
CrowdHat (Wu and Yang 2023) 68.7 296.9 0.181
STEERER (Han et al. 2023) 63.7 309.8 0.133
PET (Liu et al. 2023) 74.4 328.5 -
APGCC (Chen et al. 2024) 71.7 284.4 0.215
Gramformer (Lin et al. 2024) 72.5 316.4 0.160
mPrompt (Guo et al. 2024) 62.1 293.5 -
CLIP-EBC (Ma, Sanchez, and Guha 2024) 61.3 278.4 0.148
ZIP-B (ours) 60.1 299.0 0.104
Table 4: Comparison of our method ZIP-B with the latest approaches on the test split of NWPU-Crowd. ZIP achieves both the lowest MAE and NAE.
ZIP-B Compared with State-of-the-Art
We first compare our proposed model, ZIP-B, with recent state-of-the-art crowd counting methods on four benchmark datasets: ShanghaiTech A & B, UCF-QNRF, and the validation split of NWPU-Crowd. The results are summarized in Table 3. Our method consistently achieves the best performance across all datasets. In particular, on ShanghaiTech A, our model achieves the lowest MAE (47.81) and RMSE (75.04), outperforming the previous best result (48.8 MAE, 76.7 RMSE by APGCC). On ShanghaiTech B, we also achieve the best performance with a MAE of 5.51 and RMSE of 8.63, significantly surpassing strong baselines such as CrowdHat and STEERER. For the challenging UCF-QNRF dataset, our method sets a new state-of-the-art with 69.46 MAE and 121.88 RMSE. On NWPU-Crowd (val), we obtain a substantial improvement over prior methods, achieving 28.26 MAE and 64.84 RMSE, significantly outperforming the closest competitor, CLIP-EBC, which reports 36.6 MAE and 81.7 RMSE.
We further evaluate our method on the NWPU-Crowd test set, with results summarized in Table 4. Our ZIP-B achieves the lowest MAE (60.1) and lowest NAE (0.104) among all methods, indicating superior overall accuracy and robustness. Although the RMSE (299.0) is not the lowest, it is comparable to other top-performing methods. Notably, RMSE tends to amplify the influence of outlier errors. Given our observation of annotation noise, such as duplicate or out-of-bound head labels in the training and validation splits, RMSE may not always reflect true performance. In contrast, MAE and NAE offer more stable and representative evaluations, making our results more reliable in practice. These results clearly demonstrate the effectiveness and generalizability of our method across various crowd densities and scenes.
Comparison with Lightweight Models
ShanghaiTech A ShanghaiTech B UCF-QNRF
Method Size (M) FLOPS (G) MAE RMSE MAE RMSE MAE RMSE
LSANet (Zhu et al. 2022) 0.20 14.40 66.1 110.2 8.6 13.9 112.3 186.9
ZIP-P (ours) 0.81 6.46 71.18 109.60 8.23 12.62 96.29 161.82
ZIP-N (ours) 3.36 24.73 58.86 94.63 7.74 12.14 86.46 147.64
MobileCount (Wang et al. 2020b) 3.40 16.49 89.4 146.0 9.0 15.4 131.1 222.6
LMSFFNet-XS (Yi et al. 2023) 4.58 39.28 85.8 139.9 9.2 15.1 112.8 201.6
MobleCount×$\times 1.25$ (Wang et al. 2020b) 5.47 34.07 82.9 137.9 8.2 13.2 124.5 207.6
ZIP-T (ours) 10.53 61.39 56.36 90.55 6.67 9.90 76.02 129.40
MobleCount×$\times 2$ (Wang et al. 2020b) 13.39 63.03 81.4 133.3 8.1 12.7 117.9 207.5
ZIP-S (ours) 33.60 242.43 55.17 88.99 5.83 9.21 73.32 125.09
Table 5: Comparison of our lightweight models with state-of-the-art lightweight approaches on ShanghaiTech A & B, and UCF-QNRF. Models are grouped by parameter size: $<1$ M, $1-10$ M, $10-20$ M, $\geq 20$ M. FLOPs are measured on $1920\times 1080$ resolution. The best result in each group is highlighted, the overall best result is shown in bold.
To investigate the performance-efficiency trade-off, we compare our proposed lightweight models (ZIP-P, ZIP-N, ZIP-T, and ZIP-S) with a range of representative compact crowd counting methods. We divide models into four groups based on parameter size: (1) ultra-lightweight models ($<$1M), (2) lightweight models (1–10M), (3) mid-size models (10–20M), and (4) larger lightweight models ($\geq$20M). Table 5 presents the details.
Ultra-lightweight ($<$1M). In this group, ZIP-P achieves the best performance across all three benchmarks, with notably lower MAE on ShanghaiTech B (8.23) and UCF-QNRF (96.29) than LSANet, despite having over 50% fewer FLOPs. This demonstrates our model’s superior design in extremely constrained settings.
Lightweight (1–10M). ZIP-N significantly outperforms other methods such as MobileCount and LMSFFNet-XS across all datasets. It achieves the lowest MAE in this group on ShanghaiTech A (58.86), ShanghaiTech B (7.74), and UCF-QNRF (86.46).
Mid-size (10–20M). ZIP-T leads this group with the best results on all benchmarks. It achieves group-best scores of 56.36 MAE on ShanghaiTech A, 6.67 on ShanghaiTech B, and 76.02 on UCF-QNRF, significantly outperforming MobileCount×$\times$2 with fewer parameters and FLOPs.
Larger lightweight ($\geq$20M). ZIP-S sets a new benchmark in this group and across all lightweight models, achieving 55.17 MAE on ShanghaiTech A, 5.83 on ShanghaiTech B, and 73.32 on UCF-QNRF. These scores represent the best results among all compared lightweight models, and they also approach or surpass those of regular-sized models presented in Table 3.
Across all groups, our ZIP variants consistently outperform existing methods with comparable or fewer parameters and FLOPs. These results validate the effectiveness and scalability of our ZIP framework, making it suitable for real-world deployments under various computational constraints.
(a) Input Image.
(b) GT Density Map.
(c) Density Map (DMCount).
(d) Density Map (ZIP).
Figure 6: Qualitative comparison of DMCount and ZIP on an image from NWPU-Crowd (val). Both models utilize VGG19-based encode-decoder structure as the backbone.
Visual Comparison
We also evaluate ZIP and DMCount on the same NWPU-Crowd validation image, using an identical VGG-19 encoder–decoder backbone for fair comparison. As shown in Fig. 6, the two density maps differ markedly. DMCount produces elongated, ribbon-like streaks inside each dense block, which is a typical artifact of its transport-based loss. In contrast, ZIP yields compact, circular kernels that more closely resemble the point annotations, preserving the grid structure of the display. As for accuracy, DMCount under-counts the scene by 114 heads (-6.5%), while ZIP misses only 19 heads (-1.1%), achieving a five-fold reduction in absolute error.
Conclusion
We presented ZIP, a crowd counting framework based on zero-inflated Poisson regression. By learning a structural-zero branch to suppress background and non-head body regions, and a Poisson-rate branch to model blockwise counts, the network tackles the extreme sparsity of density maps without any explicit segmentation supervision. Comprehensive experiments show that the full model, ZIP-B, establishes new state-of-the-art MAE and RMSE on four datasets. We further introduce a family of lightweight variants (from Pico to Small) that retain these accuracy gains under tight computational budgets and consistently outperform existing models of comparable size on all four benchmarks, highlighting both the effectiveness and scalability of ZIP.
Supplementary Material
Proof of Theorem 1
For a single block with parameters $\theta\coloneqq(\pi,\lambda)$ lying in a compact set $\Theta\coloneqq[\pi_{\mathrm{min}},\pi_{\mathrm{max}}]\times[\lambda_{\mathrm{min}},\lambda_{\mathrm{max}}]\subset(0,1)\times(0,\Lambda]$  [^1] In practice, we generate $\pi$ using sigmoid and $\lambda$ using an average of finite bin centers, and thus we have $10<\pi_{\mathrm{min}}\leq\pi_{\mathrm{max}}<$ 1 and $0<\lambda_{\mathrm{min}.}\leq\lambda_{\mathrm{max}}<\infty$, define
$\mathcal{L}_{\theta}(k)$ $=-\log p_{\theta}(Y=k)$
=$=\begin{cases}-\log[\pi+(1-\pi)e^{-\lambda}],\quad&k=0,\\ -\log(1-\pi)-\lambda+k\log\lambda+\log k!,&k\geq 1.\end{cases}$
Then, for $k\in[0,\lambda_{\mathrm{max}}-1]$, we have the $L$-steep bound:
$|\mathcal{L}_{\theta}(k+1)-\mathcal{L}_{\theta}(k)|$ $=\left|\log\frac{p_{\theta}(k)}{p_{\theta}(k+1)}\right|$
$=\left|\log\frac{k+1}{\lambda}\right|$
$\leq\sup_{k,\lambda\in[\lambda_{\mathrm{min}},\lambda_{\mathrm{max}}]}\left|\log\frac{k+1}{\lambda}\right|$
$=\log\frac{k_{\mathrm{max}}+1}{\lambda_{\mathrm{min}}}\eqqcolon L,$
where $k_{\mathrm{max}}$ is the largest count that ever appears in one block in the training set. Note that $k_{\mathrm{max}}$ is restricted to a relatively small number to suppress label noise following EBC framework (Ma, Sanchez, and Guha 2024).
Following (Wang et al. 2020a), we flatten the spatial dimensions to avoid cluttered notations. For a full image with $n\coloneqq hw$ blocks, let
$g(\boldsymbol{z};\boldsymbol{\theta})=\frac{1}{n}\sum_{i=1}^{n}\mathcal{L}_{\boldsymbol{\theta}_{i}}(\boldsymbol{z}_{i}),$
For two annotation-prediction pairs $(\boldsymbol{z},\boldsymbol{\theta})$ and $(\boldsymbol{z}^{\prime},\boldsymbol{\theta})^{\prime}$:
$|g(\boldsymbol{z};\boldsymbol{\theta})-g(\boldsymbol{z}^{\prime};\boldsymbol{\theta}^{\prime})|$
$\leq$ $\frac{1}{n}\sum_{i=1}^{n}\left|\mathcal{L}_{\boldsymbol{\theta}_{i}}({\boldsymbol{z}_{i}})-\mathcal{L}_{\boldsymbol{\theta}_{i}}({\boldsymbol{z}_{i}^{\prime}})\right|+\frac{1}{n}\sum_{i=1}^{n}\left|\mathcal{L}_{\boldsymbol{\theta}_{i}}({\boldsymbol{z}_{i}^{\prime}})-\mathcal{L}_{\boldsymbol{\theta}_{i}^{\prime}}({\boldsymbol{z}_{i}^{\prime}})\right|$
$\leq$ $\frac{L}{n}\sum_{i=1}^{n}\left|\boldsymbol{z}_{i}-\boldsymbol{z}_{i}^{\prime}\right|+M\|\boldsymbol{\theta}-\boldsymbol{\theta}^{\prime}\|_{1},$
where
$M\coloneqq\sup_{k,\theta}\|\nabla_{\theta}\mathcal{L}_{\theta}(k)\|<\infty$
due to the compactness of $\Theta$.
Combining the two sums gives
$|g(\boldsymbol{z};\boldsymbol{\theta})-g(\boldsymbol{z}^{\prime};\boldsymbol{\theta}^{\prime})|$ $\leq L\|\boldsymbol{z}-\boldsymbol{z}^{\prime}\|_{1}+M\|\boldsymbol{\theta}-\boldsymbol{\theta}^{\prime}\|_{1}$
$\leq C\|(\boldsymbol{z},\boldsymbol{\theta})-(\boldsymbol{z}^{\prime},\boldsymbol{\theta}^{\prime})\|_{1}$
where $C\coloneqq\max(L,M)<\infty$. Hence, the ZIP loss is $C$-Lipschitz under the joint $\mathcal{L}_{1}$ norm.
As in the supplement of (Wang et al. 2020a), we treat the crowd counting model as $n$ identical scalar regressors. Lemma E (b) in (Wang et al. 2020a) gives
$\mathcal{R}_{\mathcal{S}}(\mathcal{L}_{\mathrm{ZIP}}\circ\mathcal{F}^{\prime})\leq C\cdot n\cdot\mathcal{R}_{\mathcal{S}}(\mathcal{H}).$
Plugging this into Theorem 26.5 in (Shalev-Shwartz and Ben-David 2014) yields, $\forall\delta\in(0,1)$, with the probability at least $1-\delta$ over the draw of the training sample $\mathcal{S}$ of size $K$,
$\mathcal{R}_{\mathcal{D}}(f_{\mathcal{S}}^{\mathrm{ZIP}},\mathcal{L}_{\mathrm{ZIP}})\leq$ $\mathcal{R}_{\mathcal{D}}(f_{\mathcal{D}},\mathcal{L}_{\mathrm{ZIP}})+2\cdot C\cdot n\cdot R_{\mathcal{S}}(\mathcal{H})+$
$5B\sqrt{\frac{2\log(8/\delta)}{K}},$
where $B=\sup_{\boldsymbol{X},\boldsymbol{z}}\mathcal{L}_{\mathrm{ZIP}}(\boldsymbol{z},f(\boldsymbol{X}))$. ∎
Inference Speed
Suffix MAE M1Pro CPU M1Pro GPU 5900X 3090
-P 8.23 9.99 118.63 122.22 1,012.01
-N 7.74 5.28 63.17 50.67 548.14
-T 6.67 1.87 27.77 18.37 249.73
-S 5.83 0.48 6.80 4.60 50.91
Table 6: Inference speed (frames per second, FPS) of the lightweight variants on ShanghaiTech B ($1024\times 768$ resolution). Results are reported on four hardware platforms.
To assess the practical deployment potential of our lightweight models, we measure the average inference speed of each ZIP variant across four representative hardware platforms: Apple M1 Pro CPU (6 performance + 2 efficiency cores), Apple M1 Pro 14-core GPU, AMD Ryzen 9 5900X CPU (12-core), and NVIDIA RTX 3090 GPU. All experiments are conducted on the ShanghaiTech B dataset using full-precision inference at an input resolution of $1024\times 768$, which reflects real-world surveillance camera scenarios. The PyTorch version is 2.7.1. Table 6 reports the average frames per second (FPS) for each model. Our smallest variant, ZIP-P, achieves the highest throughput on all platforms, reaching 122.22 FPS on AMD 5900X and over 1,000 FPS on RTX 3090, making it ideal for resource-constrained or real-time applications. On the other end, ZIP-S delivers the best accuracy (MAE = 5.83) but at a higher computational cost, with CPU-side inference limited to under 5 FPS, making it suitable for accuracy-critical but latency-tolerant use cases. ZIP-N and ZIP-T offer compelling accuracy-efficiency trade-offs. For example, ZIP-T achieves 18.37 FPS on AMD 5900X and 249.73 FPS on RTX 3090, while maintaining a competitive MAE of 6.67. These results indicate that mid-sized variants can enable near real-time performance even on high-end CPUs without requiring dedicated GPUs. Overall, our ZIP family provides a scalable solution space, allowing deployment under varying hardware and latency constraints without compromising on count accuracy.
Method ShanghaiTech A ShanghaiTech B UCF-QNRF NWPU-Crowd (Val)
MAE RMSE MAE RMSE MAE RMSE MAE RMSE
BL (Ma et al. 2019) 62.8 101.8 7.7 12.7 88.7 154.8 93.6 470.3
DMCount (Wang et al. 2020a) 59.7 95.7 7.4 11.8 85.6 148.3 70.5 357.6
EBC (Ma, Sanchez, and Guha 2024) 62.3 98.9 7.0 10.9 77.2 130.4 39.6 95.8
ZIP (ours) 58.7 93.6 6.3 10.3 73.0 126.1 38.3 90.3
Improvement over EBC 5.9% 2.9% 10.0% 5.5% 5.4% 3.2% 3.2% 5.7%
Table 7: Comparison of ZIP with other frameworks on ShanghaiTech A & B, UCF-QNRF and NWPU-Crowd. All frameworks use the VGG19-based backbone. Results demonstrate that our framework ZIP significantly outperforms other methods, achieving the lowest MAE and RMSE on all four benchmarks. Our ZIP also consistently improves the EBC framework, at most by 10.0%. The best results are highlighted in bold font, and the second best results are underscored.
Suffix Size (M) ShanghaiTech A ShanghaiTech B UCF-QNRF
MAE RMSE NAE (%) MAE RMSE NAE (%) MAE RMSE NAE (%)
-P 0.81 71.18 109.60 16.69 8.23 12.62 6.98 96.29 161.82 14.40
-N 3.36 58.86 94.63 14.15 7.74 12.14 6.33 86.46 147.64 12.60
-T 10.53 56.36 90.55 13.26 6.67 9.90 5.52 76.02 129.40 11.10
-S 33.60 55.17 88.99 11.97 5.83 9.21 4.58 73.32 125.09 10.40
-B 105.60 47.81 75.04 11.06 5.51 8.63 4.48 69.46 121.88 10.18
Table 8: Comparison of all ZIP variants on ShanghaiTech A & B, and UCF-QNRF. Best results are shown in the bold typeface and the second best are underlined.
Framework Comparison
We compare ZIP framework with several existing methods on the four benchmarks, shown in Table 7. Our approach sets a new state-of-the-art on ShanghaiTech B, UCF-QNRF and NWPU-Crowd. It also achieves the lowest MAE and the second lowest RMSE on ShanghaiTech A. Notably, when compared directly with its EBC baseline, our ZIP model demonstrates significant performance gains across all datasets, reducing the MAE by a margin of 3.2% to 10.0%. These improvements validate the efficacy of incorporating zero-inflated Poisson regression for crowd counting.
Comparison of All ZIP Variants
Table 8 reports the full results of the five ZIP variants: Pico (-P), Nano (-N), Tiny (-T), Small (-S) and Base (-B), on ShanghaiTech A, ShanghaiTech B and UCF-QNRF. The variants differ only in backbone capacity (from 0.8 M to 105 M parameters); all heads, losses and training settings are identical. From this table, we can observe three trends:
  • Monotonic accuracy gain with scale. From -P to -B every step produces lower error on all datasets. MAE on ShanghaiTech A drops from 71.2 to 47.8 (-32.8%), while UCF-QNRF MAE falls from 96.3 to 69.5 (-27.8%). A similar monotone decrease is observed for RMSE and NAE.
  • Diminishing returns. The relative improvement per additional million parameters shrinks after the -T model (10 M). For example, moving from -T to -S cuts ShanghaiTech A MAE by 3.7% whereas the jump from -N to -T yields 1.7$\times$ that reduction with fewer added parameters. This suggests that -T or -S may offer the best accuracy–efficiency trade-off for many deployments.
  • Consistent generalization across datasets. Ranking of variants is identical on ShanghaiTech A/B and UCF-QNRF, indicating that gains are not dataset-specific but stem from the scalable ZIP formulation itself.
Overall, the results demonstrate that ZIP scales gracefully: the smallest model remains competitive for edge devices, whereas the largest model achieves state-of-the-art accuracy.
Figure 7: Examples of annotation noise in NWPU-Crowd. Left column: input images with zoom-in views of the regions outlined in red. Center column: ground-truth density map with total count. Right column: density map with total count predicted by ZIP-B. Top row: mirror reflections labeled as real people. Bottom row: photo-realistic audience on a display screen left unlabeled.
Annotation Noise in NWPU-Crowd
Although NWPU-Crowd is currently the largest high-resolution benchmark, we observed non-negligible label noise. Fig. 7 presents two typical failure modes: over-labeled reflections (top row) and under-labeled screen display (bottom row).
In the top row, the red box highlights strong reflections of pedestrians on a glass facade. These mirror images have been annotated as real heads, yet they are visually less consistent with surrounding people, resulting a total count of 439. ZIP-B correctly ignores the reflections, producing a lower but more realistic count of 272.
The bottom row depicts a conference hall is shown on a large display. The audience visible inside the screen looks photo-realistic, but none of these heads are annotated (GT = 221). The model nevertheless detects them, yielding a higher prediction (671). Qualitatively, the additional detections correspond to the people shown on the screen.
Figure 8: Two typical failure cases of ZIP-B. Left: input image with a zoom-in of the critical region (red box). Center: predicted density map with total count. Right: ground-truth density map with total count.
These examples underline two points: (i) reported errors can stem from annotation artifacts rather than model failure, and (ii) the structural-zero + ZIP formulation is robust enough to down-weight unrealistic cues (mirror images) while still detecting plausible but unlabeled heads (video feed). Future benchmarks would benefit from more consistent annotation protocols or confidence tags to flag uncertain regions.
Failure Analysis
Despite its overall strong performance, ZIP-B still fails under certain visual ambiguities. Fig. 8 highlights two typical errors. The top row corresponds to false positive on man-shaped artifact: in an indoor museum scene, a post on the rope barrier strongly resembles a standing person in both shape and color, and as a result, ZIP-B misclassifies the post as a human. In the bottom row, rows of soldiers wear camouflage that blends with the background trucks. Although nearby soldiers are detected reliably, the model fails to pick up many small, low-contrast heads in the far field, leading to an under-count of -94. These failure cases suggest that incorporating additional sensing modalities, particularly infrared imagery, may help suppress person-like artifacts and recover low-contrast targets, thereby reducing both false positives and false negatives. Exploring how the ZIP formulation can be extended to multi-modal crowd counting will be an important direction for our future work.
(a) Block size = 8.
(b) Block size = 16.
(c) Block size = 32.
Figure 9: Blockwise count distributions across common crowd counting benchmarks at different block sizes (8, 16, 32). For each dataset, we first generate ground-truth pixel-level density maps from point annotations, without using Gaussian smoothing. Then, we slide a $B{\times}B$ window with stride 1 across the map, summing the density within the window to obtain its count. Bars show the percentage of windows with count values 0,1,2,3,40,1,2,3$\geq$ 40 , 1 , 2 , 3 , ≥ 4 (log-scale $y$-axis).
Blockwise Count Distribution
For each dataset’s training split, we form a pixel-level density map by placing 1 at each annotated head location and 0 elsewhere, without Gaussian smoothing being applied. We then slide an axis-aligned $B{\times}B$ window over the map with stride 1 pixel and record the integer count equal to the sum of densities in that window (equivalently, the number of annotated heads whose pixel coordinates fall inside the window). We repeat for $B\in\{8,16,32\}$ and aggregate counts over all windows from all images, converting to percentages. Fig. 9 illustrates the results. As the block size (window size) $B$ increases, zero percentages drop because each window integrates over a larger receptive field and is therefore more likely to include at least one head annotation. However, even at $B{=}32$ the distributions remain heavily zero-inflated. For instance, $>80\%$ zeros in UCF-QNRF and $>90\%$ in ShanghaiTech B and NWPU-Crowd. This indicates that annotation sparsity is a persistent regime rather than an artifact of using very fine windows. Note that increasing $B$ also coarsens supervision: spatial detail about where heads occur within a window is lost, making it harder for a crowd counting model to localize density peaks and learn fine-grained cues. In the extreme limit, $B$ equal to the entire image collapses supervision to a single global count label; the learning problem degenerates to (image-level) weakly supervised counting with no spatial guidance at all. Our ZIP formulation is motivated by operating in the practically useful regime of relatively small block size, where supervision is highly sparse yet still spatially informative, necessitating an explicit treatment of excess zeros.
Additional Qualitative Visualizations
We curate ten images from the validation split of NWPU-Crowd to demonstrate the qualitative behavior of ZIP-B in five representative scenes (two images per scene). For clarity we organize the examples into a regular group (background-only, sparse, and crowded) and a challenging group (multi-scale and low-illumination). The results are shown in Fig. 10 and Fig. 11, respectively.
Background-only (Fig. 10(a)). The structural-zero branch assigns a uniformly high zero-probability to every pixel, and the density branch outputs all-zero maps, matching the ground-truth counts of 0.
Sparse (Fig. 10(b)). With only a handful of pedestrians, the structural-zero map suppresses the background while isolating small blue blobs around true head centers; the resulting density maps reproduce the exact person counts.
Crowded (Fig. 10(c)). In the two stadium-like scenes containing several thousand people, the model still produces a clean, high-resolution density map. The total count deviates from ground truth by less than 0.5%, confirming that ZIP modeling scales to extreme densities.
Multi-scale (Fig. 11(a)). The foreground heads are large and the background heads extremely small. The structural-zero map adapts its granularity: it keeps broad foreground regions non-zero while aggressively filtering distant background pixels, enabling accurate detection across scales.
Low-illumination (Fig. 11(b)). Despite poor lighting, the structural-zero branch reliably removes dark background regions, allowing the density branch to localize heads and deliver a near-exact total count.
Together, these qualitative results show that the explicit modeling of structural zeros enables ZIP-B to generalize from empty backgrounds to dim, highly congested, and multi-scale scenes without any segmentation supervision.
(a) Background-only instances.
(b) Sparse scenarios.
(c) Highly crowded scenarios.
Figure 10: Visualized results on background-only, sparse and highly congested scenes in NWPU-Crowd (predictions by ZIP-B). Each row shows, from left to right, the input image, the predicted structural-zero map, the predicted density map, and the ground-truth density map. In the background-only examples (Fig. 10(a)), the structural-zero branch correctly classifies the entire image as background, yielding an all-zero density map. In the sparse cases (Fig. 10(b)), ZIP-B still localizes every head and reproduces the exact crowd count. As for the highly crowded scenes (Fig. 10(c)), ZIP-B can still produce accurate density maps and counts.
(a) Multi-scale scenes.
(b) Low-illumination examples.
Figure 11: Qualitative results on two challenging scenarios in NWPU-Crowd (predictions by ZIP-B): multi-scale (Fig. 11(a)) and low-illumination (Fig. 11(b)). For each row, the columns show (1) input image, (2) predicted structural-zero map, (3) predicted density map, and (4) ground-truth density map. In the multi-scale examples (Fig. 11(a)), the scenes contain a large variety of head sizes, ranging from large, close-up people to tiny, distant heads; ZIP-B suppresses background pixels while simultaneously detecting heads at all scales, yielding an accurate total count. In the low-illumination examples (Fig. 11(b)), the structural-zero branch can still filter the background, enabling the density branch to produce a clean, well-localized prediction despite poor lighting.